Aspell .33.7.1 alpha

A More Intelligent Ispell

Kevin Atkinson kevina at users sourceforge net

19th August 2001

Contents

1 Introduction 5 1.1 The Future of Aspell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.2 Comparison to other spell checker engines . . . . . . . . . . . . . . . . . . . . . . . . 6

1.2.1 Comparison to Ispell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2 Getting Started 8 2.1 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2 Obtaining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.3 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.4 Helping Out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.5 Compiling & Installing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2.5.1 Generic Install Instructions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.5.2 General Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.5.3 Curses Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.5.4 Win32 Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.5.5 Egcs 1.1 Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.6 Upgrading from version .33 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.7 Upgrading from version .32.6 . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.8 Upgrading from version .32.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.9 Upgrading from version .31.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.5.10 Upgrading from version .30 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5.11 Upgrading from version .29.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5.12 Upgrading from version .29 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5.13 Upgrading from version .28.3 . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5.14 Upgrading from version .28.2.1 . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.5.15 Upgrading from version .27.2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.5.16 Upgrading from version .25 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.5.17 Upgrading from version .24 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

3 Basic Usage 17 3.1 Spellchecking Individual Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.2 Using Aspell with other Applications . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

3.2.1 With Applications that Expect Ispell . . . . . . . . . . . . . . . . . . . . . . . 18 3.2.2 With Emacs and Xemacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.2.3 With LYX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.2.4 With VIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.2.5 With Pine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

4 Managing Word Lists 21

4.1 Creating an Individual Word List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

4.1.1 Format of the Replacement Word List . . . . . . . . . . . . . . . . . . . . . . 22 4.2 The PWLI file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.3 Dumping the contents of the word list . . . . . . . . . . . . . . . . . . . . . . . . . . 22

5 Customizing Aspell 24

5.1 Specifying Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 5.1.1 At the Command Line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 5.1.2 Via a Configuration File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.1.3 Via an Environmental Variable . . . . . . . . . . . . . . . . . . . . . . . . . . 26

5.2 The Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 5.2.1 Basic Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 5.2.2 Dictionary Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.2.3 Run-together Word Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.2.4 Filter Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.2.5 Aspell Utility Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 5.3 Dumping Configuration Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 5.4 Notes on various Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

5.4.1 Pertaining to which word lists to use . . . . . . . . . . . . . . . . . . . . . . 29 5.4.2 Notes on Various Filters and Filter Modes . . . . . . . . . . . . . . . . . . . . 31 5.4.3 Notes on the Prefix Option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 5.4.4 Notes on Typo-Analysis and the Keyboard Definition File . . . . . . . . . . . 33 5.4.5 Notes on the Different Suggestion Modes . . . . . . . . . . . . . . . . . . . . . 34

6 Writing programs to use Aspell 35

6.1 Though the Pspell API . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

6.1.1 Notes About Thread Safety . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 6.2 Through A Pipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

6.2.1 Format of the Data Stream . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 6.3 Notes of Storing Replacement Pairs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

7 Adding Support For Other Languages 39 7.1 The Language Data File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 7.2 Compiling the Word List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 7.3 Phonetic Code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

7.3.1 Syntax of the transformation array . . . . . . . . . . . . . . . . . . . . . . . . 42 7.3.2 How do I start finally? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

7.4 Controlling the Behavior of Run-together Words . . . . . . . . . . . . . . . . . . . . 46

8 How Aspell Works 49

A Changelog 50

B To Do 64 B.1 Things that will be done real soon . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 B.2 Things that need to be done . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 B.3 Things that I would like to get done . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 B.4 Things that will be done eventually . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 B.5 Good ideas that are worth implementing . . . . . . . . . . . . . . . . . . . . . . . . . 65 B.6 Things that are not likely to get implemented . . . . . . . . . . . . . . . . . . . . . . 66 B.7 Notes and Status of various items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

B.7.1 Affix Compression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 B.7.2 Extremely Large Dictionaries . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 B.7.3 General region skipping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 B.7.4 Word skipping by context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 B.7.5 Hidden Markov Model . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 B.7.6 Email the Personal Dictionary . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 B.7.7 Words With Spaces in Them . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

C Support for Gcc 2.7.2 69

D Credits 70

E Glossary and References 71

F Copyright 73 F.1 LGPL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 F.2 Ispell Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

Chapter 1

Introduction

Aspell is an Open Source spell checker designed to eventually replace Ispell. Its main feature is that it does a much better job of coming up with possible suggestions than Ispell does. In fact recent tests shows that it even does better than Microsoft Word 97’s spell checker in some cases. In addition it has both compile time and run time support for other non-English languages. Aspell is also a library however the recommend way to use aspell is through the Pspell library as the actual interface to the aspell library is constantly changing.

1.1 The Future of Aspell

Aspell .33.7 is most likely going to be the last official version of Aspell as in the near future Aspell is going to be merged into Pspell:

From: Kevin Atkinson To: aspell-announce, pspell-announce Date: 08/01/2001 Subject: Aspell and Pspell will be merged.

In the near future Aspell is going to be merged into Pspell. This will happen with the next major release of Pspell. If everything goes as planed, in the next major version of Pspell:

• Aspell will be included as part of Pspell as a module

• The Ispell module will also be included

• Most of the functionally of the Aspell command will be replaced with a new Pspell command which will work for any spell checker

• The English dictionary will no longer be included with Aspell. Instead it will be part of the Aspell-dicts package • The manuals will be switched over from LYX/L^{A}TEX to LYX/DocBook so that in can easily be

converted to other formats such as info and man.

• The current C++ interface may disappear and be replaced with a nicer C++ interface which will act as a wrapper for the C interface as I discussed earlier. If you are currently using the C++ interface and this will create a major problem for you please let me know.

This will require major reorganization of just about anything associated with Aspell and Pspell which will probably require packagers to redo there packaging scheme for Aspell and Pspell as some libraries and data files will disappear, other will be created, and the location of some of the data files will be moved. Also, as mentioned above Aspell, as you know it know, is going to disappear and be replaced with a generic utility. Some Aspell specific things, such as creating dictionaries, will remain in the Aspell utility however the executable will be renamed to aspell-util to avoid any confusion.

Also, the Aspell sourceforge project might disappear, however I am not sure about this. At very least the Aspell mailing lists and bug tracker will disappear to avoid confusion on where question and bug reports should be posted.

All of this should happen around the end of August.

1.2 Comparison to other spell checker engines

Aspell Ispell Netscape 4.0 Microsoft Word 97
Open Source x  x
Suggestion Intelligence 88-98  54  55-70?  71
Personal part of Suggestions x  x  x
Alternate Dictionaries x  x  ?  ?
International Support x  x  ?  ?

The suggestion Intelligence is based on a small test kernel of misspelled/correct word pairs. Go to http://aspell.sourceforge.net/test for more info and how you can help contribute to the test kernel. The current scores for aspell are 88 in fast mode, 93 in normal mode, and 98 in bad spellers mode see section 5.4.5 for more information about the various suggestion modes.

If you have any other information you would like to add to this chart please contact me at kevina at users sourceforge net.

1.2.1 Comparison to Ispell

1.2.1.1 Features that only Aspell has

• Does a much better job with coming up with suggestions than Ispell does or for that matter any other spell checker I have seen. If you know a spell checker that does a better job please CHAPTER 1. INTRODUCTION 7

let me know.

• Can learn from users misspellings.

• Is an actual library that others programs can link to instead of having to use it through a pipe.

• Is multiprocess intelligent. When a personal dictionary (or replacement list) is saved it will now first update the list against the dictionary on disk in case another process modified it.

• Can share the memory used in the main word list between processes.

• Support for detachable dictionaries so that more than one aspell class can use the same dictionary.

• Support for multiple personal dictionaries as well as support for special auxiliary dictionaries.

• Better support for run-together words.

• Ability to use multiple dictionaries by simply specifying it on the command line or in the configuration files.

1.2.1.2 Things that, currently, only Ispell have • Lower memory footprint

• Support for affix compression

• Perhaps better support for spell checking (La)TEX files.

• Support for spell checking Nroff files.

Getting Started

2.1 Requirements

Aspell requires gcc 2.95 (or better) as the C++ compiler. Other C++ compilers should work with some effort. Other C++ compilers for mostly POSIX compliant (Unix, Linux, BeOS, CygWin) systems should work with out any major problems provided that the compile can handle all of the advanced C++ features Aspell uses. C++ compilers for non-Unix systems might work but it will take some work. Aspell at very least requires a Unix-like environment (sh, grep, sed, tr, etc...) in order to build. Aspell also uses a few POSIX functions when necessary. Nevertheless, Aspell will compile and run using the MinGW version of gcc provided that the CygWin environment is used to to build it.

Aspell also requires the Portable Spell Checker Interface Library, otherwise known as Pspell, to be installed on your system, in the same location as Aspell will be installed in, before it will compile. Aspell requires version .12.2 or better. You can obtain the latest version of Pspell

from http://pspell.sourceforge.net/

2.2 Obtaining

The latest version can always be found at Aspell’s home page at http://aspell.sourceforge.net.

2.3 Support

Support for Aspell can be found on the Aspell mailing lists. Instructions for joining the various mailing lists (and an archive of them) can be found off the Aspell home page at http://aspell. sourceforge.net. Please use aspell-help for problems compiling and installing aspell, and aspelluser for general questions.

2.4 Helping Out

The easiest thing you can do to help out it is too send me your .aspell.≪lang≫.prepl file located in your home directory every so often. (Email kevina at users sourceforge net) The file contains data on which word pairs aspell is unable to come up with the proper suggestion. If the file does not exist is simply means that you were only using aspell in a program such as emacs which does not communicate the replacement pairs back to aspell. Sending me the file would help me improve aspell accuracy as right now I don’t have very much real data to work with.

If you are a good programmer and really want to help me consider doing one of the tasks listed in a recent post:

From: Kevin Atkinson Date: 07/18/2001 Subject: Serious Help Needed for Aspell and Pspell

In the past I have asked for help, but in specific areas and not very forcefully. Well, now I really would appreciate some help with developing Aspell and Pspell. I would like for Aspell to go to beta by the end of this summer but I really don’t see that happening unless I get some help.

Some of the areas I could use assistance in:

• Adding Affix compression support for Aspell

• Adding gettext support to Aspell and Pspell

• Making Aspell and Pspell thread safe

And lots of other smaller areas. I am looking for people who are competent programmers and have some experience with C++. I am also looking for people which have a good deal of shell programming, Perl, and autotools experience to help me clean up my build system.

Some areas such as thread safety will simply not happen until I get some help because I really do not know enough about it.

I hate to whine. However, I have put an extremely large about of unpaid time into Aspell and would appreciate some giving back from those who use Aspell in there distribution or have used the Pspell library as part of there applications, especially with those with the resources to do so such as Open Source companies.

I release that my code, especially Aspell’s, can be a bit scary with all the complex C++, however I will be glade to help any one with this part and for most of the tasks I need help with, one really don’t even need to know C++ that well. Here is a breakdown of the tasks I would like help with and the skills required.

Task: Adding Affix compression support for Aspell

Skills: Competent programmer with some C++ knowledge. I know exactly what needs to be done in this area so it will require very little knowledge of how Aspell works.

Task: Adding gettext support to Aspell and Pspell

Skills: Competent programmer with C++ knowledge and gettext knowledge. I don’t know what needs to be done in this area however they shouldn’t have to study Aspell that intensely as most of my text strings are concentrated together. They may also need to rework some of my strings into format strings (ie convert cout < "Bla.." < name < "..bla.." to something like "Bla.. $1 ..bla..") so that it can be translated correctly. However I have an idea how this should be done.

Task: Making Aspell and Pspell thread safe.

Skills: Competent programmer with good C++ and thread safety knowledge. This is the area where the C++ skills will probably be needed the most. Even though Aspell itself is not multithreaded I would like it to be thread safe so that it can be used by multi-threaded programs. I have several areas that are potently thread unsafe (such as accessing a global pool) and need some one to give me advice on how to best do this. I also have several classes which have the potential of being used by more than one thread (such as the personal dictionary) and needs to be made thread safe.

Task: Cleaning up the build system with Aspell and Pspell and similar tasks

Skills: Competent shell programmer (preferably someone with more knowledge than I have) and a good knowledge of how GNU autoconf, automake, and libtool work.

Task: Help with making the Aspell Dicts build system more robust

Skills: Competent shell programmer (preferably someone with more knowledge than I have) and a good enough knowledge of Perl to understand everything in in my "proc" script (http://aspell.

sourceforge.net/aspell-gen-0.9.tar.bz2).

Task: Miscellaneous other programming Tasks as I (or someone else) thinks of them

Skills: Competent programmer with good C++ knowledge.

Once Aspell is complete (ie gets into a beta state) you are really going to appreciate it. It will be able to do everything ispell can do and a lot more. However, I need some help in getting there.

With out some serious help it is highly likely that Aspell will not be able to be complete (ie reach a beta state) till next summer as I won’t be able to work on Aspell a great deal once School starts up again.

Thanks in advance for anyone who can offer be some help.

— Kevin Atkinson kevina at users sourceforge net

http://www.ibiblio.org/kevina/

2.5 Compiling & Installing

2.5.1 Generic Install Instructions

Before Aspell is compiled Pspell must be installed. You can obtain the latest version of Pspell from http://pspell.sourceforge.net/. Both Pspell and Aspell must have the same prefix directory in order to function correctly.

Once Pspell is installed and you have read the sections below to take care of any special requirements for you system simply type

./configure && make

or

./configure --disable-static && make

to avoid making the static libraries on a system that supports shared libraries. For additional configure options type ./configure --help. You can control what C++ compiler is used by setting the environmental variable CXX before running configure and you can control what flags are passed to the C++ compile via the environmental variable CXXFLAGS.

Aspell should then compile with out any additional user intervention. If you run into problems please first check the sections below as that might solve your problem. If it doesn’t please post a message to the aspell-help mailing list with the compiler, system you are using and any error messages that were produced. You can find more info on the aspell-help mailing list info page at

http://lists.sourceforge.net/lists/listinfo/aspell-help.

To install the program simply type

make install

And that’s all there is too it for a basic installation.

If you do not have Ispell or the traditional Unix “spell” utility installed on your system than you should also copy the compatibly scripts “ispell” and “spell” located in the scripts/ directory into your binary directory which is usually /usr/local/bin so that programs that expect ispell or spell command will work correctly.

2.5.2 General Problems

Aspell requires a specific version of Pspell. If the wrong version of Pspell is installed Aspell will not compile correctly. So before you try anything else make sure you are using the correct version of Pspell as stated in requirements (2.1) section.

Aspell does not use a released version of GNU Libtool. In previous versions of aspell this will often create problems if you inadvertently modify a file which causes Libtool to be called. However, as of Aspell .33.6.1 this should no longer be a problem and automake, autoconf, or libtool should not be called unless you specifically call them or if you configure Aspell with --enable-maintainer-mode. If you do notice any of these programs being called (and you did not configure with --enable-maintainer-mode) please let me know about it. If you have a need to modify configure.in or any of the Makefile.am’s you should install the multi-language-branch of the CVS version of libtool.

2.5.3 Curses Notes

If you are having problems compiling termios.cc than the most likely reason is due to incompatibilities with the curses implementation on your system. If this is the case than you can explicitly disable the curses library with --disable-curses. By doing this you will lose the nice full screen interface but hopefully you will be able to at least get Aspell to compile correctly.

If the curses library is installed in a non-standard location than you can specify the library and include directory with --enable-curses=≪lib≫ and --enable-curses-include=≪dir≫. Lib can either be the complete path of the library (for example “/usr/local/curses/libcurses.a”), the name of the library (for example “ncurses”) or a combined location and library in the form “-L≪lib dir≫ -l≪lib≫” (for example “-L/usr/local/ncurses/lib -lncurses”). Dir is the location of the curses header files (for example “/usr/local/ncurses/include”).

2.5.4 Win32 Notes

Aspell is now able to compile on Win32 platforms using the Win32 version of gcc. Aspell .30.1 can either be compiled with the Cgiwin or the Mingw version of Gcc 2.95 using the Cgiwin development environment. The Mingw version of Aspell will have slightly less functionality, but none of which is noticeable to the end user. In order to get the nice full screen interface with Mingw when spell checking files a curses implementation that does not require Cygwin is required. The

PDCurses (http://www.lightlink.com/hessling/PDCurses/) implementation is known to work,

other implementations may work however they have not been tested. See the previous section for information on specifying the location of the curses library and include file.

When compiling Pspell I recommend you configure with --disable-shared and --disable-ltdl. Shared libraries won’t work correctly anyway on Win32 and trying to compile the ltdl library can lead to unnecessary complications. When compiling Aspell I recommend you configure with --disable-shared. If you are planning to use Aspell outside if the Cygwin environment I strongly recommend you install Aspell in its own location (ie prefix is not /usr/local/) and compile it with --enable-win32-relocatable. Please note that Pspell and Aspell must be installed in the same location. (ie don’t install Pspell in c:/pspell and Aspell in c:/aspell. Instead install them both in c:/aspell)

If Aspell is compiled with --enable-win32-relocatable and the bindir is set to the same value as prefix (ie not ≪prefix≫/bin) then the Aspell directory (what prefix is set to) can be relocated anywhere provided that none of the data files are moved around within the Aspell directory.

The default paths for Aspell are designed for a Unix system and not a Win32 system so you might want to specify different ones when compiling Aspell. Also if the HOME environmental variable is not set Aspell will assume it is the current working directly. This may lead to your personal word lists being saved in unpredictable locations. To solve this either compile with --enable-win32-relocatable (see above) or specify the complete path of the personal and replacement word lists in aspell.conf. If Aspell is compiled with --enable-win32-relocatable than the personal word lists are saved in the prefix directory and the name is changed from “.aspell.<lang>.*” to” <lang>.*”.

2.5.5 Egcs 1.1 Notes

Aspell should now be able work with Egcs 1.1 but I have not been able to actually test it. If you have any luck one way or the other please let me know.

2.5.6 Upgrading from version .33

Even though .33.5 is a minor release it will unfortunately break binary compatibility with Aspell .33 due to the extensive changes needed to make Aspell better C++ compliant. This means applications such as Gaspell will need to be recompiled.

2.5.7 Upgrading from version .32.6

I have expanded the medium (*-med) word lists and decided to eliminate the large word lists (*-lrg) for now. However, the installing process will not automatically remove the large word lists so if you don’t want them hanging around you should delete them your self. To remove all the files remove the following files from ≪library dir≫/aspell:

american-lrg-only british-lrg-only canadian-lrg-only english-lrg-only american-lrg.multi british-lrg.multi canadian-lrg.multi

and the following files from ≪share dir≫/pspell:

en-american-lrg-aspell.pwli en-canadian-lrg-aspell.pwli en-british-lrg-aspell.pwli

2.5.8 Upgrading from version .32.1

Even though .32.5 is a minor release it breaks binary compatibly which means applications such as Gaspell will need to be recompiled.

2.5.9 Upgrading from version .31.1

The format and name of the main dictionary has changed yet again. The install process will over write the old version, so unless you are using dictionaries other than the one provided with aspell or want to have multiple versions of aspell installed you should not have to worry about this.

The apostrophe (’) is no longer considered part of the word if it appears at the end of a word. This means that you may have to manually remove words from your personal word list if you get a message similar to:

Invalid word "dogs’": The character ”’ may not appear at the end of a word.

To remove the word simply delete the line containing the word form the personal word list ( normally called “.aspell.english.pws” ).

Aspell now uses a completely new word list. This means that same words that were in the original word list may no longer appear in the current one. You may now also chose from American, British, and Canadian spelling and from two sizes medium and large. See section 5.4.1 for more information on choosing among the different choices. The original source that the word lists were created from is now found under the scowl/ directory.

2.5.10 Upgrading from version .30

The format of the main dictionary file has changed a bit. If you were able to use Aspell .30 then the old format should work. The only time the old format will NOT work is in the rare case the mmap fails. Previous version of aspell will just abort with an error when the mmap fails but the new version will attempt to read in the file using fread. Fread will fail with the old version of the main word list.

2.5.11 Upgrading from version .29.1

The format (but not the name) of the main dictionary has changed yet again. The install process will over write the old version, so unless you are using dictionaries other than the one provided with aspell or want to have multiple versions of aspell installed you should not have to worry about this.

Aspell also now depends on Portable Spell Checker Interface Library otherwise known as Pspell. Pspell must be installed before aspell will compile, you can find it at http://pspell.sourceforge.

net/.

2.5.12 Upgrading from version .29

The format and name of the main dictionary has changed which means it will need to be recompiled. The install process will remove the old files for you, so unless you are using dictionaries other than the one provided with aspell or want to have multiple versions of aspell installed you should not have to worry about this.

2.5.13 Upgrading from version .28.3

Aspell now uses namespaces which means egcs 1.0 and gcc 2.8 will no longer cut it. If this becomes a serious problem let me know as it should not be to difficult to get it working again with egcs 1.0 and gcc 2.8.

Due to the new soundslike code the the main dictionary will need to be recompiled. The build process does this automatically so unless you want to have more than one version of aspell around you should not need to worry about this.

The format and file name of the personal dictionaries has also changed. In most cases aspell will automatically detect this and convert it for you by using the following algorithm.

1. If no file exist of *.pws (for the personal word list) or *.prepl (for the personal replacement list) aspell will look for *.per or *.rpl respectfully. If that file is found it will read in the data using the old format.

2. When saving the dictionary it will save is as *.pws or *.prepl respectfully.

3. Once saved as the new format it will delete the old file.

If you have an older version of aspell around you can restore the old dictionaries by using these command

aspell.new dump personal | aspell.old create personal

aspell.new dump repl | aspell.old create repl

The new version of aspell will then leave the old files alone as long as *.pws and *.prepl exist.

Also, if the file does not end in .pws or .prepl it will try to read it in using the new format and if that fails it will read in the old format. When the file is saved it will be saved as the new format.

I am hopping I will not have to change the format of the personal dictionaries again. However, the main word list however is very likely to change in format.

2.5.14 Upgrading from version .28.2.1

The behavior of “aspell check” changed so that it will now over right the original file as creating new file was creating too many problems when used with programs like pine and vi.

2.5.15 Upgrading from version .27.2

The name of the personal word lists have changes from .aspell.per and .aspell.rpl to .aspell.≪lang≫.per and .aspell.≪lang≫.rpl respectively. ≪lang≫ is is the language name which will generally be “english”. If you wish to use your old word lists you will need to rename those files.

2.5.16 Upgrading from version .25

The format of the personal replacement dictionary has changed. So, you will either need to rename or remove the file .aspell.rpl located in your home directory. If you have information in this file you would like to preserve please send me an email.

2.5.17 Upgrading from version .24

Because the location of the main word list moved you should probably do a make uninstall (with the old version of a Aspell) before upgrading to remove the old word lists. A make uninstall will not remove any personal word lists.

Basic Usage

For a quick reference on the Aspell utility use the command “aspell --help”.

3.1 Spellchecking Individual Files

To use Aspell to spellcheck a file type in

aspell check [≪options≫] ≪filename≫ at the command line where ≪filename≫ is the file you want to check and ≪options≫ is any number of optional options. Some of the more useful ones include: --mode=≪mode> the mode to use when checking files. The available modes are none, url, email,

sgml, or tex. See section 5.4.2 for more informations on the various modes. --dont-backup don’t create a backup file. --sug-mode=≪mode≫ the suggestion mode to use where mode is one of ultra, fast, normal, or

bad-spellers. See section 5.4.5 for more information on these modes. --master=≪name≫ the main dictionary to use. The default aspell installation provided the

following dictionaries: american, british, and canadian. Please see Chapter 5 for more information on the available options. For example to check the file foo.txt:

aspell check foo.txt and to check the file foo.txt using the bad-spellers suggestion mode and the large American English dictionary:

aspell check --sug-mode=bad-spellers --master=american-lrg foo.txt

If the mode option is not given then aspell will use the extension of the file to determine the current mode. If the extension is .tex, then TEX mode will be uses, if the extension is .html, .htm, .php, or .sgml it will check the file in sgml mode, otherwise it will use url mode. Please note that the sgml-options can be used to change what extension are used for the sgml mode. See chapter 5.4.2 for more information on the various modes that can be used.

If Aspell was compiled with curses support and the TERM environmental variable is set to a capable terminal type than Aspell will use a nice full screen interface. Otherwise it will use a simpler “dumb” terminal interface where the misspelled word is surrounded by two ’*’. In either case the interface should be self explanatory.

3.2 Using Aspell with other Applications

3.2.1 With Applications that Expect Ispell

Aspell can be used as a drop in replacement for Ispell for programs that use Ispell through a pipe such as Emacs and LYX. It can also be used with programs that use simple call the ispell command and expect the original file to be overwritten with the corrected version. It support the basic features of Ispell, however it does not currently have a Nroff mode so there may be situations in which you still wish to use Ispell. Nevertheless, I have been using Aspell for Xemacs and LYX since the middle of September of 1998 with out any problems.

If you do not have Ispell installed on your system and have installed the Ispell compatibly script than you should not need to do anything as most applications will that expect Ispell will work as expected with Aspell via the Ispell compatibility script.

Otherwise, the recommended way to use Aspell as a replacement for ispell is to change the Ispell command from within the program being used. If the program uses ispell in pipe mode simple change ispell to aspell. If the program calls the ispell command to check the file change “ispell” with “aspell check”.

If that is impossible and the program uses ispell through a pipe than the run-with-aspell script can be used for programs using ispell in pipe mode. The format of the script is:

run-with-aspell ≪command≫

where ≪command≫ is the name of the program with any optional arguments.

The old method of mapping Ispell to Aspell is discouraged because it can create compatibility problems with programs that actually require Ispell such as Ispell’s own scripts.

3.2.2 With Emacs and Xemacs

The easiest way to use Aspell with Emacs or Xemacs is to add this line:

     (setq-default ispell-program-name "aspell")

to the end of your .emacs file.

For some reason version 3.0 of ispell.el (the lisp program that (x)emacs uses) want to reverse the suggestion list. To fix this add this line:

     (setq-default ispell-extra-args '("--reverse"))

after the previous line in your .emacs file and it should solve the problem.

Ispell.el, version 3.1 (December 1, 1998) and better, has the list reversing problem fixed. You can

find it at http://www.kdstevens.com/~stevens/ispell-page.html.

3.2.3 With LYX

Version 1.0 of LYX provides support for Aspell learning for users mistake feature.

To use aspell with LYX 1.0 either change the spell_command option in the lyxrc file or use the run-with-aspell utility.

3.2.4 With VIM

(The following section was written by "R. Marc", rmarc at copacetic net.)

To use aspell in vim you simply need to add the following line to your .vimrc file:

     map ˆ T :w!<CR>:!aspell check %<CR>:e! %<CR>

I use <Ctrl-T> since that’s the way you spell check in pico. In order to add a control character to your .vimrc you must type <Ctrl-v> first. In this case <Ctrl-v><Ctrl-t>.

A more useful way to use Aspell, IMHO, is in combination with newsbody (http://www.image. dk/~byrial/newsbody/) which is how I use it since vim is my editor for my mailer and my news reader.

     map ˆ T \1\2<CR>:e! %<CR>
     map \1 :w!<CR>
     map \2 :!newsbody -qs -n % -p aspell check \%f<CR>

CHAPTER 3. BASIC USAGE 20

3.2.5 With Pine

To use aspell in pine simply change the option speller to

aspell --mode=email check To change the speller option go to the main menu. Type S for setup, C for config, then W for where is. Type in speller as the word to find. The speller option should be highlighted now. Hit enter, type in the above line, and hit enter again. Then type E for exit setup and Y to save the change. If you have a strong desire to check other peoples comments change speller to

aspell check instead which will avoid switching aspell into email mode.

Managing Word Lists

4.1 Creating an Individual Word List

To create an individual main word list from a list of words use the command

aspell --lang=≪lang≫ create master ./≪base≫ < ≪wordlist≫

where ≪base≫ is the name of the word list and ≪word list≫ is the list of words separated by white space. The name of the word list will automatically be converted to all lowercase. The “./” is important because without it aspell will create the word list in the normal word list directory. If you are trying to create a word list in a language other than english check the aspell data-dir (usually /usr/share/aspell, use “aspell dump config” to find out what it is on your system) to see if a language data file exists for your language. If not you will need to create one. See chapter 7 for more information on using Aspell with other languages.

This will create the file ≪base≫ in the current directory. To use the new word list copy the file to the normal word list directory (use “aspell config” to find out what it is) and use the option - -master=≪base≫.

The compiled dictionary file is machine dependent. It is dependent on endian order, and the page size for the machine because they are mmaped in. Please do not distribute the compiled dictionaries unless you are only distributing them for a particular platform such as you would a binary. That is why is normally installed in “lib/aspell” instead of “share/aspell”.

Aspell is now also able to use special “multi” dictionaries. See section 5.4.1 form more information.

A personal and replacement word list can be created in a similar fashion.

Because Aspell does not support any sort of affix compression like Ispell does Ispell word lists will not work as is. In order to use Ispell’s word lists simply pipe the word list through “ispell -e” to expand the munched word lists.

4.1.1 Format of the Replacement Word List

The replacement word has each replacement pair on its own line in the following format

     ≪misspelled word≫: ≪correction≫

4.2 The PWLI file

In order for Aspell to be able to correctly recognize a dictionary based on the setting of the LANG environmental variable and for Pspell to be able to find your word lists each main word list installed should have at least one PWLI file associated with it in the Pspell data directory. This is normally /usr/local/share/pspell/. You can use “pspell-config pkgdatadir” to find out what it is on your system.

Each PWLI has the the following name:

≪language≫[-[≪spelling≫][-≪jargon≫]]-≪module≫.pwli

Where ≪language≫ is the two letter language code, ≪spelling≫ is the particular spelling your interested in if the languages has multiple spelling in different parts of the world such as English, ≪jargon≫ is any extra informations to distinguish the word list from other ones with the same language and spelling, and ≪module≫ is the pspell module the main word list is for.

For example:

en-aspell.pwli

en-american-aspell.pwli

en-american-medical-ispell.pwli en-american-xlg-ispell.pwli de--medical-ispell.pwli

Notice how if the spelling is left out but the jargon is not there needs to be two dashes between the language and the jargon.

Each PWLI file for an Aspell word list should then contain exactly one line which contains the full path of the main word list.

4.3 Dumping the contents of the word list

The dump command will simply dump the contents of a word list to stdout in a format than can be read back in with aspell create.

If no word list is specified the command will act on the default one. For example the command CHAPTER 4. MANAGING WORD LISTS 23

aspell dump personal will simply dump the contents of the current personal word list to stdout.

Customizing Aspell

The behavior of Aspell can be changed by any number of options which can be specified at either the command line, the environmental variable ASPELL_CONF, a personal configuration file, or a global configuration file. Options specified on the command line override options specified by the environmental variable. Options specified by the environmental variable override options specified by either of the configurations files. Finally options specified by the personal configuration file override options specified in the global configuration file. Options specified in the environmental variable ASPELL_CONF, a personal configuration file, or a global configuration file will take effect no matter how Aspell is used which includes being used by other applications.

Aspell has three basic type of options: boolean, value, and list. Boolean options are either enabled or disabled, value options take a specific value, and list options can either have entries added or removed from the list.

5.1 Specifying Options

5.1.1 At the Command Line

All options specified at the command line have the following basic format:

--≪option≫[=≪value≫]

where the ’=’ can be replaced by whitespace.

However some options also have single letter abbreviations of the form:

-≪letter≫[≪optional whitespace≫≪value≫] 5.1.1.1 Boolean

To enable a boolean option simply special the option with out any corresponding value. For example to ignore accents when checking words use “--ignore-accents”. To disable a boolean option prefix the option name with a “dont-”. For example to not ignore accents when checking words use “- -dont-ignore-accents”.

If a boolean option has a single letter abbreviation simply give the letter corresponding to either enabling or disabling the option with out any corresponding value. For example to consider runtogether words legal use “-C” or to consider them illegal use “-B”

5.1.1.2 Value

To specify a value option simply specify the option with its corresponding value. For example to set the filter mode to Tex use “--mode=tex”.

If a value option has a single letter shortcut simply specify the single letter short cut with its corresponding value. For example to use a large american dictionary use “-d american-lrg”.

5.1.1.3 List

To add a value to the list prefix the option name with a “add-” and then specify the value to add. For example to add the URL filter use “--add-filter url”. To remove a value from a list option prefix the option name with a “rem-” and then specify the value to remove. For example to remove the URL filter use “--rem-filter url”. To remove all items from a list prefix the option name with a “rem-all” without specify any value. For example to remove all filters use “--rem-all-filter”.

5.1.2 Via a Configuration File

Aspell can also accept options via a personal or global configuration file. The exact files to used are specified by the options per-conf and conf respectfully but the personal configuration file is normally “.aspell.conf” located in the HOME directory and the global one is normally “aspell.conf” which is located in the etc directory which is normally “/usr/etc” or “/usr/local/etc”. To find out the particular values for your particular system use “aspell dump config”.

Each line of the configuration file has the format:

     ≪option≫ [≪value≫]

There may any number of spaces between the option and the value however it can only be spaces, ie there is no ’=’ between the option name and the value.

Comments may also be included by preceding them with a ’#’ as anything from a “#” to a newline is ignored. Blank lines are also allowed.

Values set in the personal configuration file override those in the global file. Options specified at either the command line or via an environmental variable override those specified by either configuration file.

5.1.2.1 Boolean

To specify a boolean option simply include the option followed by a “true” to enable it or a “false” to disable it. For example to allow run-together words use “run-together true”.

5.1.2.2 Value

To specify a value option simply include the option followed by the corresponding option. For example to set the default language to german use “lang german”.

5.1.2.3 List

To add a value to the list prefix the option name with a “add-” and then specify the value to add. For example to add the URL filter use “add-filter url”. To remove a value from a list option prefix the option name with a “rem-” and then specify the value to remove. For example to remove the URL filter use “rem-filter url”. To remove all items from a list prefix the option name with a “rem-all” without specify any value. For example to remove all filters use “rem-all-filter”.

5.1.3 Via an Environmental Variable

The environmental variable ASPELL_CONF may also be used and it overrides any options set in the configuration file. The format of the string is exactly the same as the configuration file except that semicolons ( ; ) are used instead of newlines.

5.2 The Options

The following is a list of available options broken down by category. Each entry has the following format:

≪option≫[,≪single letter abbreviations≫ ] (≪type≫) ≪description≫

Where single letter options are specified as they would appear at the command line, ie with the preceding dash. Boolean single letter options are specified in the following format:

-≪abbreviation to enable≫|-≪abbreviation to disable≫

≪Option≫ is one of the following: boolean, string, file, dir, integer, or list. String, file, dir, and integer types are all value options which can only take a specific type of value.

CHAPTER 5. CUSTOMIZING ASPELL 27

5.2.1 Basic Options

conf (file) main configuration file conf-dir (dir) location of main configuration file data-dir (dir) location of language data files local-data-dir (dir) alternative location of language data files. This directory is searched before

data-dir. It defaults to the same directory the actual main word list is in (which is not necessarily dict-dir).

filter (list) add or removes a filter home-dir (dir) location for personal files ignore,-W (integer) ignore words <= n chars ignore-case (boolean) ignore case when checking words ignore-accents (boolean)ignore accents when checking words ignore-repl (boolean) ignore commands to store replacement pairs save-repl (boolean) save the replacement word list on save all lang (string) default language to use when creating a dictionary or all else fails language-tag (string) language code to use when selecting a dictionary, it follows the same format

of the LANG environmental variable on most systems. In fact it defaults to the value of the LANG environmental variable if it is set.

mode (string) sets the filter mode. Mode is one if none, url, email, sgml, or tex. (The short cut

options ’-e’ may be used for email, ’-H’ for Html/Sgml, or ’-t’ for Tex) per-conf (file) personal configuration file personal,-p (file) personal word list file name prefix (dir) prefix directory set-prefix (boolean) set the prefix based on executable location (only works on Win32 and when

compiled with --enable-win32-relocatable) repl (file) replacements list file name keyboard (file) the base name of the keyboard definition file to use (see section 5.4.4) sug-mode (mode) suggestion mode = ultra | fast | normal | bad-spellers (see section 5.4.5)

CHAPTER 5. CUSTOMIZING ASPELL 28

5.2.2 Dictionary Options

The following options may be used to control which dictionaries to use and how they behave (see section 5.4.1 for more information):

master,-d (string) base name of the main dictionary to use. The default Aspell installation pro-

vided the following dictionaries: american, british, and canadian. dict-dir (dir) location of the main word list extra-dicts (list) extra dictionaries to use strip-accents (boolean) strip accents from all words in the dictionary

5.2.3 Run-together Word Options

These may be used to control the behavior of run-together words (see section 7.4 for more information):

run-together,-C|-B (boolean) consider run-together words legal run-together-limit (integer) maximum numbers that can be strung together run-together-min (integer) minimal length of interior words

5.2.4 Filter Options

These options modify the behavior of the various filters (see section 5.4.2 for more information):

add|rem-email-quote (list) email quote characters email-margin (integer) num chars that can appear before the quote char sgml-check (list) sgml tags to always check. sgml-extension (list) sgml file extensions. tex-command (list) TEX commands tex-check-comments (boolean) check TEX comments

5.2.5 Aspell Utility Options

These options are may only be specified at the command line as there are aspell utility specific:

backup,-b|-x (boolean) create a backup file by appending “.bak” to the file name. (Only applies when the command is check)

time (boolean) time load time and suggest time in pipe mode.

reverse (boolean) reverse the order of the suggestions list.

5.3 Dumping Configuration Values

To find out the current value of all the options use the command “aspell dump config”. This will dump the current configuration to standard output. The format of the contents dumped is such that it can be used as either the global or personal configuration file.

5.4 Notes on various Options

5.4.1 Pertaining to which word lists to use

5.4.1.1 How Aspell Selects an Appropriate Dictionary

Aspell will go through the following steps to find an appropriate dictionary:

1. If the master options is set in any fashion (via the command line, the ASPELL_CONF environmental variable, or a configuration file) look for a dictionary of that name. If one could not be found complain.

2. If the language-tag (not lang) option or LANG environmental variable is set and master option is not then use it (giving preference to language-tag over LANG) to search for an appropriate dictionary. Aspell will use the same strategy that Pspell does, which is based on the installed .pwli files, to find an appropriate word list. For more information of how this is done see the Pspell manual.

3. If 2 fails than look for a dictionary of the same name of current setting of the lang options. This will currently work even if the language name is invalid, but this fact should not be relied upon as it is an implementation detail.

4. Finally, if all else fails, complain.

5.4.1.2 About Multi Dictionaries

As with precious versions of aspell you can specify the main dictionary to use via the -d or --master option. However as of Aspell .32 you can now also:

1. Specify more than word list to use with the extra-dicts option.

2. Optionally have all accents striped form the word lists using strip-accents option. This is not the same thing as the ignore-accents option. As enabling the ignore-accents would accept both cafe and café (notice the accent on the e), but only enabling strip-accents would only accent cafe, even if café is in the original dictionary. Specify strip-accents is just like using a word list with out the accents.

3. Specify special “multi” dictionaries.

A “multi” dictionary is a special file which basically a list of dictionary files to use. A multi dictionary must end is .multi and has roughly the same format of a configuration file where the two valid keys are add and strip-accents. The add key is used for adding individual word lists, or other “multi” files. The strip-accents key is used to control if accents are striped from the dictionaries. Unlike the global strip-accent option this option only effects word lists that came after the option. For example:

strip-accents yes add english strip-accents no add must-accent

will strip accents from the english word list but not the must-accent word list. If the global stripaccents option is specified the local strip-accents options are ignored.

5.4.1.3 Provided Word Lists

Aspell now provides multi dictionaries for three variates of english: american-med, british-med, and canadian-med. The word lists themselves all contain accented words however the stripaccents option is enabled by default for all the individual word lists. If you wish to use the accented words you can set the global strip-accents option to false or create a new multi word list.

Great care has been taken so that that only one spelling for any particular word is included in the main list. When two variants were considered equal I randomly picked one for inclusion in the main word list. Unfortunately this means that my choice in how to spell a word may not match your choice. If this is the case you can try to include one of the special variant dictionaries with the add-extra-dict option. You can chose from english-variant-0, english-variant-1, and englishvariant-2. Each of these word lists included all the others from the previous variant level, thus there is no need to include more than one. English-variant-0 includes most variants which are considered almost equal, english-variant-1 include variants which are also generally considered acceptable, and english-variant-2 contains variants which are seldom used. These special variant dictionaries are an experimental feature so please let me know if you take advantage of them. If no one seams to be using them I may no longer provide them in a future release of Aspell.

Many other dictionary sizes and varieties can be created. See the scowl/ directory in the source distribution for information on the different varieties you can create and section 4 for how to create an individual dictionary.

5.4.2 Notes on Various Filters and Filter Modes

Aspell now has rudimentary filter support. You can either select from individual filters or chose a filter mode. To select a filter mode use the mode option. You may chose from none, url, email, sgml, and tex. The default mode is url. Individual filters can be added with the option add-filter and remove with the rem-filter option. The currently available filters are url, email, sgml, tex as well as a bunch of filters which translate the text from one format to another.

5.4.2.1 None Mode

This mode is exactly what it says. It turns off all filters.

5.4.2.2 Url Filter/Mode

The url filter/mode skips over URL’s, host names, and email addresses. Because this filter is almost always useful and rarely does any harm it is enabled in all modes except none. To turn it off either select the none mode or use rem-filter option after the desired mode is selected.

5.4.2.3 Email Filter/Mode

The email filter/mode skips over quoted text. It currently does not support skipping over headers however a future version should. In the mean time I suggest you use Aspell with Newsbody which can

be found at http://home.worldonline.dk/~byrial/newsbody/. The option email-skip controls

the number of characters that can appear before the email quote char, the default is 10. The option add|rem-email-quote controls the characters that are considered quote characters, the default is “>’ and ’|’.

5.4.2.4 SGML Filter/Mode

The sgml filter/mode will skip over sgml commands. It currently does not handle nested < > unless they are in quotes. It also does it handle the null end tag (net) minimization feature of sgml such as

<emphasis/important/

The option add|rem-sgml-check controls which sgml tags should always be checked. The default is “alt”.

The option add|rem-sgml-extension controls which file extensions are recognized as sgml/html files. The default is html, htm, php, and sgml. The extension are not case sensitive so extensions like .HTM will also be recognized.

The sgml mode also enables a filter which will recognize sgml charter commands such as &amp; and convert it into the proper iso8859-1 character. Currently only the iso8859-1 character set is used however in future versions it will convert it to the encoding that is specified in the language date file. You can specifically turn on this filter by enable the SGML&≪charset≫/≪charset≫ filter.

5.4.2.5 TEX Filter/Mode

The tex (all lowercase) filter/mode skips over TEX commands and parameters and/or options to certain command. It also skips over TEX comments by default. The option [dont-]tex-check-comments controls whether or not aspel will skip over TEX comments. The option add|remtex-command controls which TEX commands should have certain parameters and/or options also skipped over. Commands that are not specified will have all there parameters and/or options checked. The format for each item is

≪command≫ ≪a list of p,P,o and Os≫

The first item is simple the command name. The second item controls which parameters to skip over. A ’p’ skips over a parameter while a ’P’ won’t. Similar an ’o’ will skip over an optional parameter while a ’O’ won’t. The first letter on the list will apply to the first parameter, the second letter will apply to the second parameter etc. If there are more parameters than letters Aspell will simply check them as normal. For example the option

add-tex-command rule pp

will skip over the first two parameters of the “rule” command while the option

add-tex-command foo Pop

will check the first parameter of the “foo” command, skip over the next optional parameter, if it is present, and will skip over the second parameter — even if the optional parameter is not present — and will check any additional parameters.

A’*’ at the end of the command is simply ignored. For example the option

enlargethispage p

will ignore the first parameter in both enlargethispage and enlargethispage*.

To remove a command simple use the rem-tex-command option. For example

rem-tex-command foo

will remove the command foo, if present, from the list of TEX commands.

5.4.3 Notes on the Prefix Option

The prefix option is there to allow Aspell to easily be relocated. Changing prefix will change all directory names relative to the new prefix that are not explicitly set. For example if prefix was “/usr/local/aspell” and dict-dir has a default value of “/usr/local/aspell/dict” than changing prefix to “/opt/aspell” will also change the default value of dict-dir to “/opt/aspell/dict”. Note that modifying prefix will only effect the default compiled in values of directories. If a directory option is explicitly given a value than changing the value of prefix has no effect on that directory option.

5.4.4 Notes on Typo-Analysis and the Keyboard Definition File

Aspell .33 and better will, in general, give a higher priority to certain misspelling which are likely to be due to typos such as “teh” instead of “the” or “hapoy” instead of “happy”. However in order to do this well Aspell needs to know the layout of the keyboard. The keyboard definition file simply identifies keys that are right next to each other. The file has an extension of .kbd and each line consists of two letters corresponding to two keys that are right next to each other. For example the line “as” will indicate that ’a’ and ’s’ are right next to each other. If “as” is listed as a entry it is not necessary to list “sa” as an entry as that will be done automatically. Also by “right next to each other” I mean to keys that are close enough together that it is easy to type one instead of the other. On most keyboards this means keys that are to the left or to the right of each other and not keys that are below or above it.

The default for this option is normally “standard”. However the default can be changed via the language data file. The normal default, “standard”, should work well for most QWERTY like keyboard layouts. It may need minor adjusting for foreign keyboards and will need to be completely rewritten for a Dvorak layout. When creating a keyboard definition file for a foreign language please keep in mind that Aspell completely ignores accents when scoring words so that the key ’o’ and ’ö’ will appear to be the same key to aspell even if they are in fact separate keys on your keyboard.

5.4.5 Notes on the Different Suggestion Modes

In order to understand what these suggestion modes do, a basic understanding of how aspell works is required. See section 8 for that. The suggestion modes are as follows.

ultra This method will use the fastest method available to come up with decent suggestions. This currently means that it will look for soundslikes within one edit distance apart without doing any typo analysis. It is slower than Ispell by a factor of 1.5 to 2 when a single word list is used. It speed is only minor affected by the size of the word list, if at all, but it is strongly effected by the number of word lists use. In this mode Aspell gets about 87% of the words from my small test kernel of misspelled words. (Go to http://aspell.sourceforge.net/test for more info on the test kernel as well as comparisons of this version of Aspell with previous versions and other spell checkers.)

fast This method is like ultra except that it also performs typo analysis unless it is turned off by setting the keyboard to none. The typo analysis brings words which are likely to be due to typos to the beginning of the list but slows things down by a factor of about two. This mode should get around the same number of words that the ultra method does.

normal This method looks for soundslikes within two edit distance apart and perform typo-analysis unless it is turned off. Is is around 10 times slower than fast mode with the english word list but returns better suggestions. Its speed is directly proportional to the size of the word list. This mode gets 93% of the words.

bad-spellers This method also looks for soundslikes within two edit distances apart but is more tailored for the bad speller where as fast or normal are more tailed to strike a good balance between typos and true misspellings. This mode never performs typo-analysis and returns a huge number of words for the really bad spellers who can’t seam to get the spelling anything close to what it should be. If the misspelled word looks anything like the correct spelling it is bound to be found somewhere on the list of 100 or more suggestions. This mode gets 98% of the words.

Writing programs to use Aspell

There are two main ways to use aspell from within your application. Through the Pspell API or though a pipe. The Aspell API can be used directly but that is not recommended as the actual Aspell API is constantly changing.

6.1 Though the Pspell API

To use Aspell through the Pspell API please see the Pspell manual.

6.1.1 Notes About Thread Safety

Read-only Aspell methods and functions should be thread safe as long as exceptions, new, delete, delete[], and STL allocators are thread safe. To the best of my knowledge gcc and egcs meet these requirements. It is up to the programmer to make sure multiple threads do not do thing such as change the dictionaries and add or delete items from the personal or session dictionaries.

6.2 Through A Pipe

When given the pipe or -a command aspell goes into a pipe mode that is compatible with “ispell -a”. Aspell also defines its own set of extensions to ispell pipe mode.

6.2.1 Format of the Data Stream

In this mode, Aspell prints a one-line version identification message, and then begins reading lines of input. For each input line, a single line is written to the standard output for each word checked for spelling on the line. If the word was found in the main dictionary, or your personal dictionary, then the line contains only a ’*’.

If the word is not in the dictionary, but there are suggestions, then the line contains an ’&’, a space, the misspelled word, a space, the number of near misses, the number of characters between the beginning of the line and the beginning of the misspelled word, a colon, another space, and a list of the suggestions separated by commas and spaces.

Finally, if the word does not appear in the dictionary, and there are no suggestions, then the line contains a ’#’, a space, the misspelled word, a space, and the character offset from the beginning of the line. Each sentence of text input is terminated with an additional blank line, indicating that ispell has completed processing the input line.

These output lines can be summarized as follows:

OK: *

Suggestions: & ≪original≫ ≪count≫ ≪offset≫: ≪miss≫, ≪miss≫, ...

None: # ≪original≫ ≪offset≫

When in the -a mode, Aspell will also accept lines of single words prefixed with any of ’*’, ’&’, ’@’, ’+’, ’-’, ’˜ ’, ’#’, ’!’, ’%’, or ’ˆ ’. A line starting with ’*’ tells ispell to insert the word into the user’s dictionary. A line starting with ’&’ tells ispell to insert an all-lowercase version of the word into the user’s dictionary. A line starting with ’@’ causes ispell to accept this word in the future. A line starting with ’+’, followed immediately by a valid mode will cause aspell to parse future input according the syntax of that formatter. A line consisting solely of a ’+’ will place ispell in TEX/L^{A}TEX mode (similar to the -t option) and ’-’ returns aspell to its default mode (but these commands are obsolete). A line ’˜ ’, is ignored for ispell compatibility. A line prefixed with ’#’ will cause the personal dictionaries to be saved. A line prefixed with ’!’ will turn on terse mode (see below), and a line prefixed with ’%’ will return ispell to normal (non-terse) mode. Any input following the prefix characters ’+’, ’-’, ’#’, ’!’, ’˜ ’, or ’%’ is ignored, as is any input following. To allow spell-checking of lines beginning with these characters, a line starting with ’ˆ ’ has that character removed before it is passed to the spell-checking code. It is recommended that programmatic interfaces prefix every data line with an uparrow to protect themselves against future changes in Aspell.

To summarize these:

*≪word≫ Add a word to the personal dictionary

&≪word≫ Insert the all-lowercase version of the word in the personal dictionary

@≪word≫ Accept the word, but leave it out of the dictionary

# Save the current personal dictionary

˜ Ignored for ispell compatibility.

+ Enter TEX mode.

CHAPTER 6. WRITING PROGRAMS TO USE ASPELL 37

+≪mode≫ Enter the mode specified by ≪mode≫.

- Enter the default mode. ! Enter terse mode % Exit terse mode ˆ Spell-check the rest of the line

In terse mode, Aspell will not print lines beginning with ’*’, which indicate correct words. This significantly improves running speed when the driving program is going to ignore correct words anyway. In addition to the above commands which are designed for Ispell compatibility Aspell also supports its own extension. All Aspell extensions follow the following format.

$$≪command≫ [data]

Where data may or may not be required depending on the particular command. Aspell currently supports the following command.

cs ≪option≫,≪value≫ Change a configuration option. cr ≪option≫ Prints the value of a configuration option. s ≪word1≫,≪word2≫ Returns the score of the two words based roughly on how aspell would

score them. Sw ≪word≫ Returns the soundlike equivalent of the word. Sl ≪word≫ Returns a list of words that have the same soundlike equivalent. Pw ≪word≫ Returns the phoneme equivalent of the word. pp Returns a list of all words in the current personal wordlist. ps Returns a list of all words in the current session dictionary. l Returns the current language name. ra ≪mis≫,≪cor≫ Add the word pair to the replacement dictionary for latter use. Returns nothing.

Anything returned is returned on its own line line. All lists returned have the following format

≪num of items≫: ≪item1≫, ≪item2≫, ≪etc≫

(Part of the preceding section was directly copied out of the Ispell manual)

CHAPTER 6. WRITING PROGRAMS TO USE ASPELL 38

6.3 Notes of Storing Replacement Pairs

As of version .27 of Aspell storing replacements pairs has a memory. Which means if you first store the replacement pair:

sicolagest -> psycolagest then store the replacement pair

psycolagest -> psychologist The replacement pair

sicolagest -> psychologist will also get stored so that you don’t have to worry about it.

Adding Support For Other Languages

Before you consider adding support for Aspell for your language first check if it is already done via the Aspell-Dicts project. As of the release of support for the following languages is now available:

Breton  [br]
Catalan  [ca]
Czech  [cs]
Danish  [da]
Dutch  [nl]
Esperanto [eo]
Faroese  [fo]
French  [fr] (Standard [fr_FR] and Swiss French [fr_CH] and in three sizes: small, medium and large)
German  [de] (Standard [de_DE] and Swiss German [de_CH])
Italian  [it]
Norwegian [no]
Polish  [pl]
Portuguese [pt] (Standard [pt_PT] and Brazilian [pt_BR])
Russian  [ru]
Spanish  [es]
Swedish  [sv]

You can find all of these packages at the aspell home page (http://aspell.sourceforge.net/).

If your language is not listed above please send me a note and I will work with you on adding support. The instructions below still apply however for this version of Aspell however they may not once the Aspell is merged into Pspell (see section 1.1).

Adding a language to aspell is fairly straightforward. You need to at very least create the language data file, and compile a new word list. You should also create a PWLI file for each of your word lists so that your new word lists will work correctly with the LANG environmental variable and with Pspell.

Please note, however, that Aspell international support is not 100% done yet. More information on my future planes for international support in aspell can be found at http://aspell.sourceforge.

net/international/.

7.1 The Language Data File

The basic format of the language data data is the same as it for aspell configuration file. It is named ≪lang≫.dat and is located in the architecture independent data dir for aspell (option datadir) which is usually ≪prefix≫/share/aspell. Use “aspell config” to find out where it is in your installation.

The language data file has several mandatory fields, and several optional ones. All fields are case sensitive and should be in all lower case.

The two mandatory fields are name and charset. Name is the name of the language and should be the same as the file name (without the .dat). Charset is the charset aspell will expect the word lists to be formatted in. You may chose from any of the iso-8859-* character sets as well as, koi8-f, koi8-r, and viscii. If your language can fit in the plain old ASCII character set use iso8859-1. If you use some other character set for your language other than the ones listed here drop me a note and I will look into adding support for it.

The optional fields are special, soundslike, keyboard and a bunch of options to specify how runtogether words are handles. Special is for non letter characters that can appear in your language such as the ’ and -. The format for the value is a list separated by spaces. Each item of the list has the following format

≪char≫ ≪begin≫≪middle≫≪end≫

≪char≫ is the non letter character in question. ≪begin≫,≪middle≫,≪end≫ are either a ’-’ or a ’*’. A star for ≪begin≫ means that the character at the beginning of the word, a ’-’ means it can’t. The same is true for ≪middle≫ and ≪end≫. For example the entry for the ’ in english is:

     ' -*-

To include more than one middle character just list them one after another on the same line. for example to make both the ’ and the - a middle character use the following line in the language data file:

     special ' -*- - -*-

The soundslike option, if present, should be the name of the soundslike data for the language. The data is expected to be in the file ≪name≫_phonetic.dat.

If the name is generic a really generic soundslike algorithm will be used which consists of striping all the vowels and removing all accents. I recommend first using the generic algorithm and then, after aspell is working with the new language, work on the transformation array.

If the soundslike name is none then no soundslike lookup table will be used. This will reduce the size of the compiled word list by around 50% but at the sacrifice of suggestion quality. If the soundslike is none than the soundslike for the word will simply be the word itself in lowercase, will all accents stripped. For languages with phonetic spelling the difference will not be very noticeable. However, for languages with non-phonetic spelling there will be a noticeable difference. The difference you notice will depend on the quality of the soundslike data file. If you do not notice much of a difference for a language with non-phonetic spelling that is a good indication that the soundslike data is not rough enough—or the words you are trying are not that badly misspelled.

The keyboard option specifies the base name of the keyboard definition file to use. See section 5.4.4 for more information.

The options to control how run-together words are handled are the same as the are in the normal configurations files. Please see section 7.4 for more information.

7.2 Compiling the Word List

Once you have a working language data file installed in the right place you are ready to compile the main word list. See section 4 to find out what to do. This section also includes instructions for creating the PWLI file.

7.3 Phonetic Code

(The following section was written by Björn Jacke, bjoern.jacke at gmx de)

Aspell is in fact the spell checker that comes up with the best suggestions if it finds an unknown word. One reason is that it does not just compare the word with other words in the dictionary (like Ispell does) but also uses phonetic comparisons with other words.

The new table driven phonetic code is very flexible and setting up phonetic transformation rules for other languages is not difficult but there can be a number of stumbling stones — that’s why I wrote this section.

The main phonetic code is free of any language specific code and should be powerful enough to allow setting up rules for any language. Anything which is language specific is kept in a plain text file and can easily be edited. So it’s even possible to write phonetic transformation rules if you don’t have any programming skills. All you need to know is how words of the language are written and how they are pronounced.

7.3.1 Syntax of the transformation array

In the translation array there are two strings on each line; the first one is the search string (or switch name) and the second one is the replacement string (or switch parameter). The line

     version ≪version≫

is also required to appear somewhere in the translation array. The version string can be anything but it should be changed when ever the a new version of the translation array is released. This is important because it will keep Aspell from using a compiled dictionary with the wrong set of rules. For example if when coming up with suggestion for “hallo” aspell will use the new rules to come up with the soundslike say “H*L*” but if hello is stored in the dictionary using the old rules as “HL” instead of “H*L*” aspell will never be able to come up with hello. So to solve this problem Aspell checks if the version strings match and abort with an error if they don’t. Thus it is important to update it when ever a new version of the translation array is releases. This is only a problem with the main word list as the personal word lists are now stored as simple word lists with a single header line (ie, no soundslike data).

Each non switch line represents one replacement (transformation) rule. Words beginning with the same letter must be grouped together; the order inside this group does not depend on alphabetical issues but it gives priorities; the higher the rule the higher the priority. That’s why the first rule that matches is applied. In the following example:

     GH     _
     G      K

“GH → _” has higher priority than “G → K”. “_” represents the empty string “”. If “GH → _” would stand after “G → K”, the second rule would never match because the algorithm would stop searching for more rules after the first match. The above rules transform any “GH” to an empty string (delete them) and transform any other “G” to “K”.

At the end of the first string of a line (the search string) there may optionally stand a number of characters in brackets. One (only one!) of these characters must fit. It’s comparable with the [ ] brackets in regular expressions. The rule “DG(EIY) → J” for example would match any “DGE”, “DGI” and “DGY” and replace them with “J”. This way you can reduce several rules to one.

Behind the search string there can stand one or more dashes (-). Those search strings will be matched totally but only the beginning of the string will be replaced. Furthermore for these rules no follow-up rule will be searched (what this is will be explained later). The rule “TCH-- → _” will match any word containing “TCH” (like “match”) but will only replace the first character “T” with an empty string. The number of dashes determines how many characters from the end will not be replaced. After the replacement the search for transformation rules continues with the not replaced “CH”!

If a “<” is appended to the search string, the search for replacement rules will continue with the replacement string and not with the next character of the word. The rule “PH< → F” for example would replace “PH” with “F” and then again start to search for a replacement rule for “F… ”. If there would also be rules like “FO → O” and “F → _” then words like “PHOXYZ” would be transformed to “OXYZ” and any occurrences of “PH” that are not followed by an “O” will be deleted like “PHIXYZ → IXYZ”. The second replacement however is not applied if the priority of this rule is lower than the priority of the first rule.

Priorities are added to a rule by putting a number between 0 and 9 at the end of the search string, for example “ING6 → N”. The higher the number the higher is the priority.

Priorities are especially important for the previously mentioned follow-up rules. Follow-up rules are searched beginning from the last string of the first search string. This is a bit complicated but I hope this example will make it more clear:

     CHS      X
     CH       G
     HAU--1   H
     SCH      SH

In this example “CHS’ in the word “FUCHS” would be transformed to “X”. If we take the word “DURCHSCHNITT” the things look a bit different. Here “CH” belongs together and “SCH” belongs together and both are spoken separately. The algorithm however first finds the string “CHS” which may not be transformed like in the previous word “FUCHS”. At this point the algorithm can find a follow up rule. It takes the last character of the first matching rule (“CHS”) which is “S” and looks for the next match, beginning from this character. What it finds is clear: It finds “SCH → SH”, which has the same priority (no priority means standard priority, which is 5). If the priority is the same or higher the follow-up rule will be applied. Let’s take a look at the word “SCHAUKEL”. In this word “SCH” belongs together and may not be torn apart. After the algorithm has found "SCH → SH" it searches for a follow-up rule for "H"+"AUKEL". It finds "HAU--1 → H", but does not apply it because its priority is lower than the one of the first rule. You see that this is a very powerful feature but it also can easily lead to mistakes. If you really don’t need this feature you can turn it off by putting the line

     followup      0

at the beginning of the phonetic table file. As mentioned, for rules containing a ‘-’ no follow-up rules are searched but giving such rules a priority is not totally senseless because they self can be follow-up rules and in that case the priority makes sense again. Follow-up rules of follow-up rules are not searched because this is in fact not needed very often.

The control character "^{∧}" says that the search string only matches at the beginning of words so that the rule "RH^{∧}→ R" will only apply to words like "RHESUS" but not "PERHAPS". You can append another "^{∧}" to the search string. In that case the algorithm treats the rest of the word totally separately from first matched string in at beginning. This is useful for prefixes whose pronunciation does not depend on the rest of the word and vice versa like "OVER^{∧}^{∧}" in English for example.

The same way as "^{∧}" works does "$" only apply on words that end with the search string. "GN$ → N" only matches on words like "SIGN" but not "SIGNUM". If you use "^{∧}" and "$" together, both of them must fit "ENOUGH^{∧}$ → NF" will only match the word "ENOUGH" and nothing else.

Of course you can combine all of the mentioned control characters but they must occur in this order: < - priority ˆ $. All characters must be written in CAPITAL letters.

If absolutely no rule can be found — might happen if you use strange characters for which you don’t have any replacement rule — the next character will simply be skipped and the search for replacement rules will continue with the rest of the word.

If you want double letters to be reduced to one you must set up a rule like "LL- → L". If double letters in the resulting phonetic word should be allowed, you must place the line

     collapse_result     0

at the beginning of your transformation table file; otherwise set the value to ‘1’. The English rules for example strip all vowels from words and so the word "GOGO" would be transformed to "K" and not to "KK" (as desired) if collapse_result is set to 1. That’s why the English rules have

collapse_result set to 0.

7.3.2 How do I start finally?

Before you start to write an array of transformation rules, you should be aware that you have to do some work to make sure that things you do will result in correct transformation rules.

7.3.2.1 Things that come in handy

First of all you need to have a large word list of the language you want to make phonetics for. It should contain about as many words as the dictionary of the spell checker. If you don’t have such a list, you will probably find an Ispell dictionary at http://fmg-www.cs.ucla.edu/geoff/ ispell-dictionaries.html which will help you. You can then make affix expansion via ispell -e and then pipe it trough \tr " " "\n" to put one word on each line. After that you eventually have to convert special characters like ‘é’ from Ispell’s internal representation to latin1 encoding. sed s/e'/é/g for example would replace all e’ with é.

The second is that you know how to use regular expressions and know how to use grep. You should for example know that

     grep ˆ [ˆ aeiou]qu[io] wordlist | less

will show you all words that begin with any character but a, e, i, o or u and then continue with ‘qui’ or ‘quo’. This stuff is important for example to find out if a phonetic replacement rule you want to set up is valid for all words which match the expression you want to replace. Taking a look at the regex(7) man page is a good idea.

7.3.2.2 What the phonetic code should do

Normal text comparison works well as long as the typer misspells a word because he pressed one key he didn’t really want to press. In this cases mostly one character differs from the original word.

In cases where the writer didn’t know about the correct spelling of the word however the word may have several characters that differ from the original word but usually the word would still sound like the original word. Someone might think for example that ‘tough’ is spelled ‘taff’. No spell checker without phonetic code will come to the idea that this might be ‘tough’ but a spell checker who knows that ‘taff’ would be pronounced like ‘tough’ will make good suggestions to the user. Another example could be ‘funetik’ and ‘phonetic’.

From this examples you can see that the phonetic transformation should not be too fussy and too precise. If you implement a whole phonetic dictionary as you can find it in books this will not be very useful because then there could still be many characters differing from the misspelled and the desired word. What you should do if you implement the phonetic transformation table is to reduce the number of used letters to the only really necessary ones.

Characters that sound similar should be reduced to one. In English language for example ‘Z’ sounds like ‘S’ and that’s why the transformation rule “Z → S” is present in the replacement table. ‘PH’ is spoken like ‘F’ and so we have a “PH → F” rule.

If you take a closer look you will even see that vowels sound very similar in English language: ‘contradiction’, ‘cuntradiction’, ‘cantradiction’ or ‘centradiction’ in fact sound nearly the same, don’t they? Therefore the English phonetic replacement rules not only reduce all vowels to one but even remove them all (removing is done by just setting up no rule for those letters). The phonetic code of ‘contradiction’ is ‘KNTRTKXN’ and if you try to read this letter-monster loud you will hear that it still sound a bit like ‘contradiction’. You also see that ‘D’ is transformed to ‘T’ because they nearly sound the same.

If you think you have found a regularity you should always take your word list and grep for the corresponding regular expression you want to make a transformation rule for. An example: If you come to the idea that all English words ending on ‘ough’ sound like ‘AF’ at the end because you think of ‘enough’ and ‘tough’. If you then grep for the corresponding regular expression by “grep -i ough$ wordlist” you will see that the rule you wanted to set up is not correct because the rule doesn’t fit to words like ‘although’ or ‘bough’. So you have to define your rule more precisely or you have to set up exceptions if the number of words that differ from the desired rule is not so big.

Don’t forget about follow-up rules which can help in many cases but which also can lead to many confusions and side effects. It’s also important to write exceptions in front of the more general rules ("GH" before "G" etc.).

If you think you have set up a number of rules that may produce some good results try them out! If you run Aspell as “aspell --lang=≪your language≫ pipe” you get a prompt at which you can type in words. If you just type words Aspell checks them and eventually makes suggestions if they are misspelled. If you type in “$$Sw ≪word≫” you will see the phonetic transformation and you can test out if your work does what you want.

Another good way to control if changes you apply to your rules don’t have any evil side effects is to create another list from your word list which contains not only the word of the word list but also the corresponding phonetic version of this word on the same line. If you do this one time before the change and one time after the change you can make a diff (see man diff) to see what really changed. To do this use the command “aspell --lang=≪your language≫ soundslike”. In this mode aspell will output the the original word and then its soundslike separated by a tab character for each word you give it. If you are interested in seeing how the algorithm works you can download a set of useful

programs from http://members.xoom.com/maccy/spell/phonet-utils.tar.gz. This includes a

program that produces a list as mentioned above and another program which illustrates how the algorithm works. It uses the same transformation table as Aspell and so it helps a lot during the process of creating a phonetic transformation table for Aspell.

During your work you should write down your basic ideas so that other people are able to understand what you did (and you still know about it after a few weeks). The English table has a huge documentation appended for example.

Now you can start experimenting with all the things you just read and perhaps set up a nice phonetic transformation table for your language to help Aspell to come up with the best correction suggestions ever seen also for your language. Take a look at the Aspell homepage to see if there is already a transformation table for your language. If there is one you might also take a look at it to see if it could be improved.

If you think that this section helped you or if you think that this is just a waste of time you can send any feedback to bjoern.jacke@gmx.de.

7.4 Controlling the Behavior of Run-together Words

Aspell has support for either unconditionally accepting run-together words or only accepting certain words in compound formation.

Support for unconditionally accepting run-together words can either be turned on in the language data file or as a normal option via the run-together option. The run-together-limit options controls the maximum number of words that can be strung together, the default is normally 255. The run-together-min options controls the minimal length the individual components of the run together word can be, the default is normally 3. Both the run-together-limit and run-together-min option may be specified in both the language data file or as a normal. The run-together-mid option, which may only be specified in the language data file, may be used to specify up to three optional characters that may appear between individual words.

In order for aspell to conditionally only accept certain words in compounds those words must be flagged when the compiled word list is being created. The format for each entry is

≪word≫:C[1][2][3]≪middle char≫

The 1, 2, and 3 control if the word is allowed to appear in the begging, middle, or end of the compound, respectfully. More than one position flag may be specified. If none of them are specified it as assumed that the word may appear anywhere. The C is optional if 1, 2, or 3 is specified. The ≪middle char≫ represents an optional character that may appear after the word in the formation of the compound if the word is not at the end of the compound. If the letter is lowercase than the character may appear after the word, if it is in uppercase then that letter must appear after the compound. Only one letter may be specified and it must also be in the list of middle letters specified via the run-together-mid option. The run-together-limit option may also be used to specify the maximum number of words to string together.

For example the word list:

beg:1 mid:2 end:3 any:C never must:CM maybe:Cm

Means that the word “beg” may only appear at the begging of a word, the word “mid” at the middle, the word “end” at the end, and the word “any” any place. The word “never” is never accepted in a compound unless the run-together option is set. The word “must” may appear anywhere however it must be followed by an “m”, while the word maybe may be followed by an “m”. Given the above word list the following compounds or legal:

begmidend begany mustmend maybeend maybemend

are all legal, but the following are not:

begmid mustend neverany Individual words such as “beg” are always accepted.

When the run-together option is not set Aspell will only accept words that have been flagged in a run-together word. When the run-together option is set aspell will accept words which are as least as long as the value specified in the run-together-min option. If the words length is less than run-together-min then it will only accept the word if it has been flagged. When the run-together option is not set the run-together-min option is ignored all together.

Currently Aspell only supports run-together words when checking if a word is in the dictionary. When coming up with suggestions Aspell treats the word as a normal word and does not do anything special. This means that the suggestions will be virtually meaningless when the actual word is a run-together. I plan on more intelligently supporting run-together words when coming up with suggestions in a future version of Aspell.

How Aspell Works

The magic behind my spell checker comes from merging Lawrence Philips excellent metaphone algorithm and Ispell’s near miss strategy which is inserting a space or hyphen, interchanging two adjacent letters, changing one letter, deleting a letter, or adding a letter.

The process goes something like this.

1. Convert the misspelled word to its soundslike equivalent (its metaphone for English words).

2. Find all words that have a soundslike within one or two edit distances from the original words soundslike. The edit distance is the total number of deletions, insertions, exchanges, or adjacent swaps needed to make one string equivalent to the other. When set to only look for soundslikes within one edit distance it tries all possible soundslike combinations and check if each one is in the dictionary. When set to find all soundslike within two edit distance it scans through the entire dictionary and quickly scores each soundslike. The scoring is quick because it will give up if the two soundslikes are more than two edit distances apart.

3. Find misspelled words that have a correctly spelled replacement by the same criteria of step number 2 and 3. That is the misspelled word in the word pair (such as teh -> the) would appear in the suggestions list as if it was a correct spelling.

4. Score the result list and return the words with the lowest score. The score is roughly the weighed average of the weighed edit distance of the word to the misspelled word and the soundslike equivalent of the two words. The weighted edit distance is like the edit distance except that the various edits have weights attached to them.

5. Replace the misspelled words that have correctly spelled replacements with their replacements and remove any duplicates that might arise because of this.

Please note that the soundslike equivalent is a rough approximation of how the words sounds. It is not the phoneme of the word by any means. For more details about exactly how each step is performed please see the file suggest.cc. For more information on the metaphone algorithm please see the data file english_phonet.dat.

Appendix A

Changelog

Changes from .33.7 to .33.7.1 (Aug 20, 2001)

• Minor manual fixes.

• Compile fix for Gcc 3.0 and Solaris.

Changes from .33.6.3 to .33.7 (Aug 2, 2001)

• Updates to Autoconf 2.50 and switched to the HEAD branch of libtools.

• Fixed a bug which caused Aspell to crash when typo-analysis is not used such as when sugmode is fast or bad spellers.

• Added support for typo-analysis even when a soundslike is not used.

• Fixed a bug which causes extended charters to display incorrectly on some platforms

• Compile fixes so that it will compile with Gcc 3.0.

• Compile fixed which should allow Aspell to compile with Egcs 1.1. I have not been able to actually test it though. Please let me know at kevina@users.sourceforge.net if you have have tried with Egcs 1.1.

• Compile and configuration script fixes so that USE_FILE_INO will properly be defined and Aspell will compile correctly when it is defined.

• More ANSI C++ compliance fixes.

APPENDIX A. CHANGELOG 51

Changes from .33.6.2 to .33.6.3 (June 3, 2001)

• Fixed a build problem in the manual/ directory by including manual-text and manual-html in the distribution.

Changes from .33.6.1 to .33.6.2 (June 3, 2001)

• Compile fix so that Aspell will work correctly when not installed in /usr/local.

• Avoided regenerating the manual unless configured with enable-maintainer-mode.

• Added the missing documentation files in the scowl directory.

Changes from .33.6 to .33.6.1 (May 29, 2001)

• Fixed a formating problem with the manual involving ≪.

• Added a note about creating PWLI files.

• Removed the space after between the -L and the directory name in the pspell-module/ Makefile which caused problems on some platforms.

• Added the configure option AM_MAINTAINER_MODE to avoid enabling rules which often causes generated build files to be rebuild with the wrong version of Libtool by default. I don’t know why I didn’t think to do this a long time ago.

Changes from .33.5 to .33.6 (May 18, 2001)

• Fixed a minor bug where some words would have random compound tags attached to them.

• Fixed a compile problem on many platforms where fileno is defined as a macro.

• Updated the description for a few of Aspell’s options.

• Removed the note of Aspell not being able to run when compiled with the upcoming Gcc 3.0 compiler as things seam to work now.

• Added a note about Aspell not being able to compile with Egcs 1.1.

• Added hack to deal with Libtool’s interdependencies problem. See bug #416981 for Pspell for more info.

APPENDIX A. CHANGELOG 52

Changes from .33 to .33.5 (April 5, 2001)

• “dump master” correctly detects which dictionary and language to use based on the LANG environmental variable.

• Fixed a problem on Win32 which involves path names that begin with ≪Drive Letter≫:.

• Bug fixes and enhancements so that Aspell can once again run under MinGW. You can even use the new full screen interface if Aspell is compiled with PDCurses.

• Some major modifications to make Aspell more C++ compliant in order to get Aspell to compile under the upcoming Gcc 3.0 compiler. This included only using STL features found in the standard version of C++. (Which means Aspell will no longer require using the SGI version of the STL) This should also make compiling C++ under non-gcc compilers a lot simpler. Please not that Aspell still has some problems with the upcoming Gcc 3.0 compiler (see section ?? for more info).

• Minor changes to remove some -Wall warnings.

• Added a hack to that Aspell will properly compile as a shared library under Solaris.

• Added a few import missing words to the English word list.

Changes from .32.6 to .33 (January 28, 2001)

• Added a new new curses based interface to replace the dumb terminal interface everyone has been bitching about.

• Added the ability to give higher priority to words such as "the" instead of "teh" which are likely to be due to typos.

• Reorganized the manual so that it is hopefully easier to follow.

• Ability to automatically select the best dictionary to used based on the setting of the LANG environmental variable.

• Expanded the medium dictionary size to include more words which included the original words found in ispell and eliminated the large size for now.

• Added three special variant add-on dictionaries.

• Switched to the multi-language branch of the CVS version of libtool.

APPENDIX A. CHANGELOG 53

Changes from .32.5 to .32.6 (Nov 8, 2000)

• Fixed a bug that where Aspell will crash when reading in accented characters on some platforms. This fixes bug # 112435.

• Fixed some other bugs so that it will run under Win32 under CygWin. Unfortunately it still won’t run properly under Mingw.

• Fixed the mmap test in configure so that it won’t fail on some platforms that use munmap(char *, int) instead of munmap(void *, int).

• Upgraded to the latest CVS version of libtool which fixed the problem with using GNU Make under Solaris.

• Added an option to copy files instead of using symbolic links for the special “multi” dictionary files.

Changes from .32.1 to .32.5 (August 18, 2000)

• Changed my email from kevinatk at home com to kevina at users sourceforge net please make a note of the new email address.

• Added an option to control if the personal replacement dictionary is saved when the save_all_wls method is called.

• Brought back the ability to dump the master word list even in the case of the special “multi” lists.

• Added a large number of hacker related words and some other slang terms to the medium size word list.

• Added an “ispell” and “spell” compatibility script for systems which don’t have ispell installed. They are located in the scripts/ directory and are not installed by default.

• Manual fixes.

• Added a note on not using GNU Make on Solaris.

Changes from .32 to .32.1 (August 5, 2000)

• Minor compile fixes for recent gcc snapshot.

• Fixed naming of pwli files.

• Fixed a bug when aspell will crash when used with certain single letter flags. This bug was most noticeable when used with Emacs.

• Word list changes, see SCOWL Readme.

• Other miscellaneous changes.

Changes from .31.1 to .32 (July 23, 2000)

• Added support for optionally doing without the soundslike data.

• Greatly reduced the amount of memory used when creating word lists.

• Added support for ignoring accents when coming up with suggestions.

• Added support for local-data-dir which is searched before data-dir.

• Added support for specifying which words may be used in compounds and where they may be used.

• Added support for having more than one main word list as well as a special “multi” word list files which will allow multiple word lists to be treated as one.

• Aspell now uses a completely new word list.

• The apostrophe (’) is no longer considered part of the word when it as at the end of the word such as in “ dogs’ ”.

Changes from .31 to .31.1 (June 18, 2000)

• Fixed a bug where Aspell would not create a complete dictionary file on some platforms when the data is 8-bit.

• Added a workaround so Aspell will work with ispell.el 3.3.

• Minor compile fixes so it would compile better with the very latest gcc (CVS Version).

• Removed note about compiling in Win32 as I was now able to get it to work.

Changes from .30.1 to .31 (June 11, 2000)

• Added support for spell checking run together words.

• Added an option to produce a list of misspelled words from standard input.

• More robust error reporting when reading in language data files.

APPENDIX A. CHANGELOG 55

• Fixed a bug when that will cause Aspell to crash if the “special” line was not defined in the language data file.

• Update Pspell Module.

• Minor bug fixes.

• Added cross references in “The Aspell utility Chapter” for easier use.

Changes from .30 to .30.1 (April 29, 2000)

• Ported Aspell to Win32 platforms.

• Portability fixes which may help aspell compile on other platforms.

• Aspell will no longer fail if for some reason the mmap fails, instead it will just read the file in as normal and free the memory when done.

• Minor changes in the format of the main word list as a result of the changes, the old format should still work in most cases.

• Fixed a bug when aspell was ignoring the extension of file names such as .html or .tex when checking files.

• Fixed a bug when aspell will go into an infinite loop when creating the main word list from a word list which has duplicates in it.

• Minor changes to the manual for better clarity.

Changes from .29.1 to .30 (April 2, 2000)

• Fixed many of the capitalization bugs found in previous versions of Aspell.

• Changed the format of the main word list yet again.

• Fixed a bug so that “aspell check” will work on the PowerPC.

• Added ability to change configuration options in the middle of a session.

• Added words from /usr/dict/words found on most Linux systems as well as a bunch of commonly used abbreviation to the word list.

• Fixed a bug when aspell will dump core after reporting certain errors when compiled with gcc 2.95 or higher. This involved reworked the Exception heritage to get around a bug in gcc 2.95.

• Added a few more commands to the list of default commands the TEX filter knows about.

APPENDIX A. CHANGELOG 56

• Aspell will now check if a word only contains valid characters before adding it to any dictionaries. This might mean that you have to manually delete a few words from your personal word list.

• Added option to ignore case when checking a document.

• Adjusted the parameters of the “normal” suggest mode to so that significantly less far fetched results are returned in cases such as tomatoe, which went from 100 suggestions down to 32, at the expense of getting slightly lower results (less than 1%),

• Improved the edit distance algorithm for slightly faster results.

• Removed the $$m command in pipe mode, you should now use “$$cs mode,≪mode≫” to set the mode and “$$cr mode” to find out the current mode.

• Reworked parts of Aspell to use Pspell services to avoid duplicating code.

• Added a module for the newly released Pspell. It will get installed with the rest of aspell.

• Miscellaneous other bug fixes.

Changes from .29 to .29.1 (Feb 18, 2000)

• Improved the TEX filter so that it will accept ’@’ at the begging of a command name and ignore trailing ’*’s. It also now has better defaults for which parameters to skip.

• Reworked the main dictionary so that it can be memory mapped in. This decreases startup time and allows multiple aspell processes to use the same memory for the main word list. This also also made Aspell 64 bit clean so that it should work on an alpha now.

• Fix so that aspell can compile on platforms that gnu as is not available for.

• Fixed issue with flock so it would compile on FreeBSD.

• Minor changes in the code to make it more C++ compliant although I am sure there will still be problems when using some other compiler other than gcc or egcs.

• Added some comments to the header files to better document a few of the classes.

Changes from .28.3 to .29 (Feb 6, 2000)

• Fixed a bug in the pipe mode with lines that start with “ˆ $$”.

• Added support for ignoring all words less than or equal to a specified length • New soundslike code based thanks to the contribution of Björn Jacke. It now gets all of its

data from a table making it easier for other people to add soundslike code for their native language. He also converted the metaphone algorithm to table form, eliminating the need for the old metaphone code.

• Major redesign of the suggestion code for better results.

• Changed the format of the personal word lists. In most cases it should be converted automatically.

• Changed the format of the main word list.

• Name space cleanup for more consistent naming. I now use name spaces which means that gcc 2.8.* and egcs 1.0.* will no longer cut it.

• Used file locks when reading and saving the personal dictionaries so that it truly multiprocess safe.

• Added rudimentary filter support.

• Reworked the configuration system once again. However, the changes to the end user who does not directly use my library should be minimal.

• Rewrote my code that handles parsing command line parameters so that it no longer uses popt as it was causing to many problems and didn’t integrate well with my new configuration system.

• Fixed pipe mode so that it will properly ignore lines starting with ’˜ ’ for better ispell compatibility.

• Aspell now has a new home page at http://aspell.sourceforge.net/. Please make note of the new URL.

• Miscellaneous manual fixes and clarifications.

Changes from .28.2.1 to .28.3 (Nov 20, 1999)

• Fixed a bug that caused aspell to crash when spell checking words over 60 characters long.

• Reworked “aspell check” so that

1. You no longer have to hit enter when making a choice.

2. It will now overwrite the original file instead of creating a new file. An optional backup can be made by using the -b option.

• Fixed a few bugs in data.cc.

APPENDIX A. CHANGELOG 58

Changes from .28.2 to .28.2.1 (Aug 25, 1999)

• Fixed the version number for the shared library.

• Fixed a problem with undefined references when linking to the shared library.

Changes from .28.1 to .28.2 (Aug 25, 1999)

• Fixed a bunch of bugs in the language and configuration classes.

• Minor changed in the code so that it could compile with the new gcc 2.95(.1).

• Changed the output of “dump config” so that default values are given the value "<default>". This means that the output can be used to created a configuration file.

• Added notes on using aspell with VIM.

Changes from .28 to .28.1 (July 27, 1999)

• Removed some debug output

• Changed notes on compiling with gcc 2.8.* as I managed to get it to compile on my school account

• Avoided included stdexcept in const_string.hh so that I could get to compile on my schools account with gcc 2.8.1.

Changes from .27.2 to .28 (July 25, 1999)

• Provided an iterator for the replacement classes.

• Added support for dumping and creating and merging the personal and replacement word lists.

• Changed the aspell utility command line a bit, it now used popt.

• Totally reworked aspell configuration system. Now aspell can get configuration from any of 5 sources: the command line, the environmental variable ASPELL_CONF, the personal configuration file, the global configuration file, and finally the compiled in defaults.

• Totally reworked the language class in preparation for my new language code. See http:

  //aspell.sourceforge.net/international/ for more information of what I have in store.

APPENDIX A. CHANGELOG 59

• Added some options to the configure script: --enable-dict-dir=DIR, --enable-doc-dir=DIR, - -enable-debug, and --enable-opt

• Removed some old header files.

• Reorganized the directory structure a bit

• Made the text version of the manual pages slightly easier to read

• Used the \url command for urls for better formating of the printed version.

Changes from .27.1 to .27.2 (Mar 1, 1999)

• Fixed a major bug that caused aspell to dump core when used without any arguments

• Fixed another major bug that caused aspell to do nothing when used in interactive mode.

• Added an option to exit in Aspell’s interactive mode.

• Removed some old documentation files from the distribution.

• Minor changes on to the section on using Aspell with egcs.

• Minor changes to remove -Wall warnings.

Changes from .27 to .27.1 (Feb 24, 1999)

• Fixed a minor compile problem.

• Updated the section on using Aspell with egcs to it it more clear why the patch is necessary.

Changes from .26.2 to .27 (Feb 22, 1999)

• Totally reworked the C++ library which means you may need to change some things in your code.

• Added support for detachable and multiple personal dictionaries in the C++ class library.

• The C++ class library now throws exceptions.

• Reworked aspell ability to learn from users misspellings a bit so that it now has a memory. See section 6.3 for more information.

• Upgraded autoconf to version 2.13 and automake to version 1.4 for better portability.

APPENDIX A. CHANGELOG 60

• Fixed the configuration so the “make dist” will work. For now on aspell will be distributed with “make dist”.

• Added support to skip over URL’s, email addresses and host names.

• Added support for dumping the master and personal word list. You can now also merge a personal word list. Type aspell -help for help on using this feature.

• Reorganized the source code.

• Started using proper version numbers for the shared library.

• Fixed a bug that caused a spell to crash when adding certain replacement pairs.

• Fixed the problem with duplicate lines when exiting pipe mode for good.

Changed from .26.1 to .26.2 (Jan 3, 1998)

• Fixed another compile problem. Hopefully this time it will really compile OK on other peoples machines.

Changed from .26 to .26.1 (Jan 3, 1998)

• Fixed a small compile problem in as_data.cc.

Changed from .25.1 to .26 (Jan 3, 1999)

• Fixed a bug that causes duplicates items to be displayed in the suggestion list for good. (If it still does it please send be email.)

• Added the ability for aspell to learn form the users misspellings.

• Library Interface changes. Still more to come....

• Is now multiprocess safe. When a personal dictionary (or replacement list) is saved it will now first update he list against the dictionary on disk in case another process modified it.

• Fixed the bug that caused duplicate output when used non interactively in pipe mode.

• Dropped support for gcc 2.7.2 as the C++ compiler.

• Updated the How Aspell Works (8) Chapter.

• Added support for the ASPELL_DATA_DIR environmental variable.

APPENDIX A. CHANGELOG 61

Changes from .25 to .25.1 (Dec 10, 1998)

• Fixed the version number so that Aspell reports the correct version number.

• Changed the note on gcc 2.7.2 compilers to make it clear that only the C++ compiler can not be gcc 2.7.2, it is ok if the C compiler is gcc 2.7.2.

• Updated the TODO list and reorganized it a bit.

• Fixed the directory so that all the documentation will get installed in ${prefix}/doc/aspell instead of half of it in ${prefix}/doc/aspell and half of it in ${prefix}/doc/kspell.

Changes from .24 to .25 (Nov 23, 1998)

• Total rework of how the main word list is stored. Start up time decreased to about 1/3 of what it was in .24 and memory usage decreased to about 2/3. (When used with the provided word list on a Linux system). Also the format and default locations of the main word list data files changed in the process and the data is now machine dependent. The personal word list format, however, stayed the same.

• Changed the scoring method to produce slightly better results with words like the vs. teh. And other simpler misspellings where two letters are swapped.

• Fixed the very unpredictable behavior of the ’*’, ’&’, ’@’ commands in the pipe mode.

• Added documentations for Aspell pipe mode (also known as ispell -a compatibility mode)

• Added a bunch of Aspell specific extensions to the pipe mode and documented them.

• Documented the to_soundslike and soundslike methods for the aspell class.

• Changed the scoring method to produce better results for words like “fone” vs “phone” and other words that have a spelling that doesn’t directly relate to how the word sounds by using the phoneme equivalent of the word in the scoring of it.

• Added the to_phoneme and have_phoneme methods to the SC_Language class.

• Added the to_phoneme method to the aspell class.

• Added the framework for being able to learn from the users misspelling. Right now it just keep a log of replacements.

• Re did stl_rope-30.diff. For some reason the version of patch on my system refused it.

• Rewrite of the “Using as a replacement for Ispell” section and added the run-with-aspell utility as a replacement of the old method of mapping Ispell to Aspell.

• Fixed a bug that caused duplicate words to appear in the suggestion list.

APPENDIX A. CHANGELOG 62

Changes from .23 to .24 (Nov 8, 1998)

• Fixed my code so that it can once again compile with g++ 2.7.2.

• Rewrote the How It Works chapter.

• Rewrote the Requirement section and added noted on compiling with g++ 2.7.2.

• Added a To Do chapter.

• Added a Glossary and References chapter.

• Other minor documentation improvements.

• Internal code documentation improvements.

Changes from .22.1 to .23 (Oct 31, 1998)

• Minor documentation fixes.

• Changed the scoring strategy for words with 3 or less letters. This cut the number of words returned for these roughly in half.

• Expanded the word list to also include american.0 and american.1 from the Ispell distribution. It now includes english.0, english.1, american.0 and american.1 from the directory languages/english provided with Ispell 3.1.20.

• Added a link to the location of the latest Ispell.el in the documentation.

• Started a C interface and added some rough documentation for it.

Changes from .22 to .22.1 (Oct 27, 1998)

• Minor bug fixes. I was deleting arrays with delete rather than delete[]. I was suppressed that this had not created a problem.

• Added a simple test program to test for a memory leak present on some systems. (Only debian slink at the moment.) See the file memleak-test.cc for more info.

Changes from .21 to .22 (Oct 26, 1998)

• Major redesign or the scoring method. It now uses absolute distances rather than relative scores for more consistent results. See suggest.cc for more info.

APPENDIX A. CHANGELOG 63

• Suggest code rewritten is several places however the core process stayed the same.

• The suggest_ultra method temporally does nothing. It should be working again by the next release.

Changes from .20 to .21 (Oct 13, 1998)

• Added documentation for aspell::Error

• Changed the library name from libspell to libaspell. It should never have been libspell in the first place. Sorry for the incompatibility.

• Added as_error.hh to the list of files copied to the include directory so that you can actually use the library outside of the source dir.

• Fixed bug that caused a segmentation fault with words where the only suggestions was inserting a space or hyphen such as in “ledgerline”.

• Added the score method to aspell.

• Changed the scoring method to deal with word when the user uses "f" in place of "ph" a lot better.

Changes from .11 to .20 (Oct 10, 1998)

• Name change. Everything that was Kspell is now Aspell. Sorry, the name Kspell was already used by KDE and I didn’t want to cause any confusion.

• Fixed a bug that causes a segmentation fault when the HOME environmental variable doesn’t exist.

Changes from .10 to .11 (Sep 12, 1998)

• Overhaul of the SC_Language class

• Added documentation for international support

• Added documentation for the C++ library

• Other minor bug fixes.

Appendix B

To Do

Words in bold indicate how you should refer to the item when discussing it with me or others.

B.1 Things that will be done real soon

These items should get done within a release or two.

• Totally rewrite the aspell international support. See http://aspell.sourceforge.net/ international/ for more information.

• Rework the aspell check function to provide support for using any number of filters which will be needed for international support.

• Add support for more intelligently coming up with suggestions for words that are run-togethers.

B.2 Things that need to be done

Things items will eventually be implemented as I know they are important however I am not sure when they will get done.

• Figure out a way for Aspell to work better with extremely large dictionaries.

B.3 Things that I would like to get done

These items will eventually be implemented. I hope to have them all done before I move aspell to beta testing. They are in the approximate order of when they will get done.

B.4 Things that will be done eventually

I plan on doing these things eventually. It is just a matter of getting around to it.

B.5 Good ideas that are worth implementing

These items all sound like good ideas however I am not sure when I will get to implementing then if ever. If you are looking for a way to contribute picking up on one of these ideas would be a great way to start. They are presented in no particular order.

• Use Lawrence Philips’ new Double Metaphone algorithm. See http://aspell.sourceforge.

  net/metaphone/.

• Add support for affix compression.

• Come up with a plug-in for gEdit the gnome text editor.

• Change languages (and thus dictionaries) based on the information in the actual document.

• Come up with a nroff mode for spell checking.

• Come up with a mode that will skip words based on the symbols that (almost) always surround the word. (Word skipping by context)

• Create two server modes for Aspell. One that uses the DICT protocol and one that uses ispell -a method of communication of some arbitrary port.

• Come up with thread safe personal dictionaries.

• Use the Hidden Markov Model to base the suggestions on not only the word itself but on the context around the word.

• Having a way to email the personal dictionary and/or replacement list to a particular address either periodical or when it grows to a certain size.

• Be able to accept words with spaces in them as many languages have words such as as a word in a foreign phrases which only make sense when followed by other words.

The following good ideas where found in the ispell WISHES file so I thought I would pass them on.

• Ispell should be smart enough to ignore hyphenation signs, such as the TEX \- hyphenation indicator.

• (Jeff Edmonds) The personal dictionary should be able to remove certain words from the master dictionary, so that obscure words like "wether" wouldn’t mask favorite typos.

• (Jeff Edmonds) It would be wonderful if ispell could correct inserted spaces such as "th e" for "the" or even "can not" for "cannot".

• Since ispell has dictionaries available to it, it is conceivable that it could automatically determine the language of a particular file by choosing the dictionary that produced the fewest spelling errors on the first few lines.

B.6 Things that are not likely to get implemented

Theses ideas are not likely to get implemented any time soon.

• (None Yet)

B.7 Notes and Status of various items

B.7.1 Affix Compression

Due to the current way my spell checker works implementing affix compression would be next to impossible. Nevertheless, I do realize that for some languages affix compression is very important.

So to solve this dilemma I plan on having two different modes of my spell checker: One with affix compression that does not use soundslike pairs (much like ispell) and one without affix compression that does use soundslike.

I plan to extract the affix manipulation code from Ispell with the help of an Ispell author. The tricky part would be getting this to getting this all to work properly at tun time bases on the dictionary used.

B.7.2 Extremely Large Dictionaries

This problem extends back to the fact of the way words are index is Aspell. This problem will get resolved when I implanted the affix compression mode as only one index would be used.

B.7.3 General region skipping

I want to implement this give other people an idea of how it should be done and because I am really sick of having to spell check through url and email address.

B.7.4 Word skipping by context

This was posted on the Aspell mailing list on January 1, 1999:

I had an idea on a great general way to determine if a word should be skipped. Determine the words to skip based on the symbols that (almost) always surround the word.

For example when asked to check the following C++ code:

     cout << "My age is: " << num << endl;
     cout << "Next year I will be " << num + 1 << endl;

cout, num, and endl will all be skipped. "cout" will be skipped because it is always preceded by a <<. "num" will be skipped because it is always preceded by a <<. And "endl" will be skipped because it is always between a << and a ;.

Given the following html code.

     <table width=50% cellspacing=0 cellpadding=1>
     <tr><td>One<td>Two<td>Three
     <tr><td>1<td>2<td>3
     </table>
     <table cellspacing=0 cellpadding=1>
     </table>

table, width cellspacing, cellpadding, tr, td will all be skipped because they are always enclosed in "<>". Now of course table and width would be marked as correct anyway however there is no harm in skipping them.

So I was wondering if anyone on this list has any experience in writing this sort of context recognition code or could give me some pointers in the right direction.

This sort of word skipping will be very powerful if done right. I imagine that it could replace specific spell checker modes for Tex, Nroff, SGML etc because it will automatically be able to figure out where it should skip words. It could also probably do a very good job on programming languages code.

If you are interested in helping be out with this or just have general comments about the idea please let me know.

B.7.5 Hidden Markov Model

Knud Haugaard Sørensen suggested this one. From his email on the Aspell mailing list: consider this examples. a fone number. -> a phone number. a fone dress. -> a fine dress. the example illustrates that the right correction might depend on the context of the word. So I suggest that you take a look on HMM to solve this problem. This might also provide a good base to include grammar correction in aspell.

see this link http://www.cse.ogi.edu/CSLU/HLTsurvey/ch1node7.html

I think it is a great idea. However unfortunately it will probably be very complicated to implement. Perhaps in the far future.

B.7.6 Email the Personal Dictionary

Some one suggest in a personal email:

Have you thought of adding a function to aspell, that - when the personal dictionary has grown significantly - sends the user’s personal dictionary to the maintainer of the corresponding aspell dictionary? (if the user allows it) It would be a very useful service to the dictionary maintainers, and I think most users can see their benefit in it too.

And I replied:

Yes I have considered something like that but not for the personal dictionaries but rather the replacement word list in order to get better test data for http://aspell. sourceforge.net/test/. The problem is I don’t know of a good way to do this sense Aspell can also be used as a library. It also is not a real high priority, especially sense I would first need to learn how to send email within a C++ program.

B.7.7 Words With Spaces in Them

While this is something I would like to do it is not a simple task. The basic problem is that when tokenizing a string there is no good way to keep phrases together. So the solution is to some how add special conditions to certain words which will dictate which words can come before/after it. Then there is also a problem of how to come up with intelligent suggestions. What further complicates things is that many applications send words to Aspell a word at a time. So even if Aspell did support such a thing many applications that would use Aspell will not. So, in order for this to work applications will need to send text to Aspell a document or at least a sentence at a time. Unfortunately the framework for doing this is not there yet. It will be once I finish the filter interface. Another possible is to provide call back functions in which Aspell will be able to request the previous or next word on request. Yet again the framework for doing this is not there. Perhaps sometime in the near future.

Appendix C

Support for Gcc 2.7.2

(and other non-standard compliance compilers)

My original plan was to program in such a way that it would Aspell would compile under Gcc 2.7.2 however after releasing a rather nasty bug in 2.7.2 with nested types I have desired to drop all support for Gcc 2.7.2. As of Aspell .27 all hope for being able to compile under Gcc 2.7.2 is lost as I am now using many modern C++ features which are simply not present in Gcc 2.7.2, most notably template specialization and template members. Egcs 1.1.1 is a very good standards compliant compiler and that is now the officially supported compiler. However as Egcs 1.1.1 is relatively new and, except for namespaces, provides little new functionality I will continue to support Egcs 1.0.3. Gcc 2.8.1 should in theory work however it is so buggy I have yet to get it to compile with it nor has anyone else that I know of.

Yes, my code could be rewritten so that is could compile under Gcc 2.7.2 but Why? Using modern C++ has probably accelerated the development of this library my at least 50%. And for that matter why stop at Gcc 2.7.2 why not go all out and totally rewrite my code pure C. I hope you see my point.

However that does not mean I want to sacrifice portability unnecessarily. If you see any of part of my code that in not Standard C++ please let me know. My hope is that my code could compile on all Standard compliance C++ compilers with the addition of a few extra header files from SGI’s STL.

As a side note I think that Mozilla’s C++ portability guide (http://www.mozilla.org/docs/ tplist/catBuild/portable-cpp.html) could be summed up in one sentence: Program in the dark ages of C++.

Appendix D

Credits

• To the many authors of Ispell (including R. E. Gorin, Pace Willisson, and Geoff Kuenning) for providing me with a good word list as well as giving me a few good ideas.

• Alan Beale for going well beyond the call of duty with helping me create a better word list.

• Lawrence Philips for coming up with the original metaphone algorithm and Michael Kuhn for writing C code for the algorithm.

• Björn Jacke for coming up with a generic soundslike algorithm which gets all of its data from a file, thus eliminating almost all need for language specific code from aspell.

• To the authors of SGI STL version 3.0 and up for proving a great set of generic container classes with cut the development time of this program in half at least.

• To the LYX development team for giving me a great tool for the development of this manual.

Appendix E

Glossary and References

affix in grammar, a word element that, when added to a word, modifies its meaning or function;

prefix, infix, or suffix. Debian A 100% Open Source Linux distribution http://www.debian.org DICT protocol A TCP transaction based query/response protocol that allows a client to access

dictionary definitions from a set of natural language dictionary databases. http://www.dict.

     org/

Gnome A project to build a complete, user-friendly desktop based entirely on free software. http:

     //www.gnome.org/

GTK+ A library for creating graphical user interfaces for the X Window System. It is designed

to be small, efficient, and flexible. http://www.gtk.org/ GUI Graphics User Interface

Ispell.el Emacs interface for ispell. http://www.kdstevens.com/~stevens/ispell-page.html

Ispell An international spell checker which is just about the only decent Open Source spell checker

     out there. (except Aspell of course). http://fmg-www.cs.ucla.edu/geoff/ispell.html

Ispell -a An Ispell mode that is designed to be used by other applications though a pipe. KDE A powerful graphical desktop environment for Unix workstations. http://www.kde.org Linux A Open Source version of Unix which runs on many platforms. http://www.linux.org/ LYX An What You See is What You Mean document editor for the X environment, http://www.

     lyx.org

Open Source Software where the source code is available for anyone to extend or modify. http:

     //www.opensource.org/

APPENDIX E. GLOSSARY AND REFERENCES 72

Red Hat A commercial Linux distribution. http://www.redhat.com RPM Red Hat’s packing format also used by other Linux distributions. STL Standard Temple Library. A C++ library of container classes, algorithms, and iterators. SGI STL Silicon Graphics implantation of the STL. http://www.sgi.com/Technology/STL/ STLPort A port of SGI STL designed to run on compilers that don’t support all the latest features

     of C++. http://corp.metabyte.com/~fbp/stl/

Appendix F

Copyright

The library and utility program is Copyrighted 2000 by Kevin Atkinson under the LGPL. Certain parts of the library, as indicated at the top of the source file, are under a weaker license. The provided word list is extracted from the word lists provided with Ispell thus it is distributed under the same terms of Ispell.

The two licenses follow:

F.1 LGPL

GNU LIBRARY GENERAL PUBLIC LICENSE Version 2, June 1991

Copyright (C) 1991 Free Software Foundation, Inc.

59 Temple Place, Suite 330, Boston, MA 02111-1307 USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

[This is the first released version of the library GPL. It is numbered 2 because it goes with version 2 of the ordinary GPL.]

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public Licenses are intended to guarantee your freedom to share and change free software — to make sure the software is free for all its users.

This license, the Library General Public License, applies to some specially designated Free Software Foundation software, and to any other libraries whose authors decide to use it. You can use it for your libraries, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the library, or if you modify it.

For example, if you distribute copies of the library, whether gratis or for a fee, you must give the recipients all the rights that we gave you. You must make sure that they, too, receive or can get the source code. If you link a program with the library, you must provide complete object files to the recipients so that they can relink them with the library, after making changes to the library and recompiling it. And you must show them these terms so they know their rights.

Our method of protecting your rights has two steps: (1) copyright the library, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the library.

Also, for each distributor’s protection, we want to make certain that everyone understands that there is no warranty for this free library. If the library is modified by someone else and passed on, we want its recipients to know that what they have is not the original version, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that companies distributing free software will individually obtain patent licenses, thus in effect transforming the program into proprietary software. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

Most GNU software, including some libraries, is covered by the ordinary GNU General Public License, which was designed for utility programs. This license, the GNU Library General Public License, applies to certain designated libraries. This license is quite different from the ordinary one; be sure to read it in full, and don’t assume that anything in it is the same as in the ordinary license.

The reason we have a separate public license for some libraries is that they blur the distinction we usually make between modifying or adding to a program and simply using it. Linking a program with a library, without changing the library, is in some sense simply using the library, and is analogous to running a utility program or application program. However, in a textual and legal sense, the linked executable is a combined work, a derivative of the original library, and the ordinary General Public License treats it as such.

Because of this blurred distinction, using the ordinary General Public License for libraries did not effectively promote software sharing, because most developers did not use the libraries. We concluded that weaker conditions might promote sharing better.

However, unrestricted linking of non-free programs would deprive the users of those programs of all benefit from the free status of the libraries themselves. This Library General Public License is intended to permit developers of non-free programs to use free libraries, while preserving your freedom as a user of such programs to change the free libraries that are incorporated in them. (We have not seen how to achieve this as regards changes in header files, but we have achieved it as regards changes in the actual functions of the Library.) The hope is that this will lead to faster development of free libraries.

The precise terms and conditions for copying, distribution and modification follow. Pay close attention to the difference between a "work based on the library" and a "work that uses the library". The former contains code derived from the library, while the latter only works together with the library.

Note that it is possible for a library to be covered by the ordinary General Public License rather than by this special one.

GNU LIBRARY GENERAL PUBLIC LICENSE TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License Agreement applies to any software library which contains a notice placed by the copyright holder or other authorized party saying it may be distributed under the terms of this Library General Public License (also called "this License"). Each licensee is addressed as "you".

A "library" means a collection of software functions and/or data prepared so as to be conveniently linked with application programs (which use some of those functions and data) to form executables.

The "Library", below, refers to any such software library or work which has been distributed under these terms. A "work based on the Library" means either the Library or any derivative work under copyright law: that is to say, a work containing the Library or a portion of it, either verbatim or with modifications and/or translated straightforwardly into another language. (Hereinafter, translation is included without limitation in the term "modification".)

"Source code" for a work means the preferred form of the work for making modifications to it. For a library, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the library.

Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running a program using the Library is not restricted, and output from such a program is covered only if its contents constitute a work based on the Library (independent of the use of the Library in a tool for writing it). Whether that is true depends on what the Library does and what the program that uses the Library does. 1. You may copy and distribute verbatim copies of the Library’s complete source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and distribute a copy of this License along with the Library.

You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Library or any portion of it, thus forming a work based on the Library, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a) The modified work must itself be a software library.

b) You must cause the files modified to carry prominent notices stating that you changed the files and the date of any change.

c) You must cause the whole of the work to be licensed at no charge to all third parties under the terms of this License.

d) If a facility in the modified Library refers to a function or a table of data to be supplied by an application program that uses the facility, other than as an argument passed when the facility is invoked, then you must make a good faith effort to ensure that, in the event an application does not supply such function or table, the facility still operates, and performs whatever part of its purpose remains meaningful.

(For example, a function in a library to compute square roots has a purpose that is entirely welldefined independent of the application. Therefore, Subsection 2d requires that any applicationsupplied function or table used by this function must be optional: if the application does not supply it, the square root function must still compute square roots.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Library, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Library, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it.

Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Library.

In addition, mere aggregation of another work not based on the Library with the Library (or with a work based on the Library) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may opt to apply the terms of the ordinary GNU General Public License instead of this License to a given copy of the Library. To do this, you must alter all the notices that refer to this License, so that they refer to the ordinary GNU General Public License, version 2, instead of to this License. (If a newer version than version 2 of the ordinary GNU General Public License has appeared, then you can specify that version instead if you wish.) Do not make any other change in these notices.

Once this change is made in a given copy, it is irreversible for that copy, so the ordinary GNU General Public License applies to all subsequent copies and derivative works made from that copy.

This option is useful when you wish to copy part of the code of the Library into a program that is not a library.

4. You may copy and distribute the Library (or a portion or derivative of it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you accompany it with the complete corresponding machine-readable source code, which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange.

If distribution of object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place satisfies the requirement to distribute the source code, even though third parties are not compelled to copy the source along with the object code.

5. A program that contains no derivative of any portion of the Library, but is designed to work with the Library by being compiled or linked with it, is called a "work that uses the Library". Such a work, in isolation, is not a derivative work of the Library, and therefore falls outside the scope of this License.

However, linking a "work that uses the Library" with the Library creates an executable that is a derivative of the Library (because it contains portions of the Library), rather than a "work that uses the library". The executable is therefore covered by this License. Section 6 states terms for distribution of such executables.

When a "work that uses the Library" uses material from a header file that is part of the Library, the object code for the work may be a derivative work of the Library even though the source code is not. Whether this is true is especially significant if the work can be linked without the Library, or if the work is itself a library. The threshold for this to be true is not precisely defined by law.

If such an object file uses only numerical parameters, data structure layouts and accessors, and small macros and small inline functions (ten lines or less in length), then the use of the object file is unrestricted, regardless of whether it is legally a derivative work. (Executables containing this object code plus portions of the Library will still fall under Section 6.)

Otherwise, if the work is a derivative of the Library, you may distribute the object code for the work under the terms of Section 6. Any executables containing that work also fall under Section 6, whether or not they are linked directly with the Library itself.

6. As an exception to the Sections above, you may also compile or link a "work that uses the Library" with the Library to produce a work containing portions of the Library, and distribute that work under terms of your choice, provided that the terms permit modification of the work for the customer’s own use and reverse engineering for debugging such modifications.

You must give prominent notice with each copy of the work that the Library is used in it and that the Library and its use are covered by this License. You must supply a copy of this License. If the work during execution displays copyright notices, you must include the copyright notice for the Library among them, as well as a reference directing the user to the copy of this License. Also, you must do one of these things:

a) Accompany the work with the complete corresponding machine-readable source code for the Library including whatever changes were used in the work (which must be distributed under Sections 1 and 2 above); and, if the work is an executable linked with the Library, with the complete machine-readable "work that uses the Library", as object code and/or source code, so that the user can modify the Library and then relink to produce a modified executable containing the modified Library. (It is understood that the user who changes the contents of definitions files in the Library will not necessarily be able to recompile the application to use the modified definitions.) b) Accompany the work with a written offer, valid for at least three years, to give the same user the materials specified in Subsection 6a, above, for a charge no more than the cost of performing this distribution.

c) If distribution of the work is made by offering access to copy from a designated place, offer equivalent access to copy the above specified materials from the same place.

d) Verify that the user has already received a copy of these materials or that you have already sent this user a copy.

For an executable, the required form of the "work that uses the Library" must include any data and utility programs needed for reproducing the executable from it. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable.

It may happen that this requirement contradicts the license restrictions of other proprietary libraries that do not normally accompany the operating system. Such a contradiction means you cannot use both them and the Library together in an executable that you distribute.

7. You may place library facilities that are a work based on the Library side-by-side in a single library together with other library facilities not covered by this License, and distribute such a combined library, provided that the separate distribution of the work based on the Library and of the other library facilities is otherwise permitted, and provided that you do these two things:

a) Accompany the combined library with a copy of the same work based on the Library, uncombined with any other library facilities. This must be distributed under the terms of the Sections above.

b) Give prominent notice with the combined library of the fact that part of it is a work based on the Library, and explaining where to find the accompanying uncombined form of the same work.

8. You may not copy, modify, sublicense, link with, or distribute the Library except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, link with, or distribute the Library is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

9. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Library or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Library (or any work based on the Library), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Library or works based on it.

10. Each time you redistribute the Library (or any work based on the Library), the recipient automatically receives a license from the original licensor to copy, distribute, link with or modify the Library subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

11. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Library at all. For example, if a patent license would not permit royalty-free redistribution of the Library by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Library.

If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply, and the section as a whole is intended to apply in other circumstances.

It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice.

This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

12. If the distribution and/or use of the Library is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Library under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

13. The Free Software Foundation may publish revised and/or new versions of the Library General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

Each version is given a distinguishing version number. If the Library specifies a version number of this License which applies to it and "any later version", you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Library does not specify a license version number, you may choose any version ever published by the Free Software Foundation.

14. If you wish to incorporate parts of the Library into other free programs whose distribution conditions are incompatible with these, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

15. BECAUSE THE LIBRARY IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE LIBRARY, TO THE EXTENT PERMITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE LIBRARY "AS IS" WITHOUT WARRANTY OF ANY KIND, EI-THER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE LIBRARY IS WITH YOU. SHOULD THE LIBRARY PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

16. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRIT-ING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE LIBRARY AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUEN-TIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE LIBRARY (IN-CLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INAC-CURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE LIBRARY TO OPERATE WITH ANY OTHER SOFTWARE), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS

How to Apply These Terms to Your New Libraries

If you develop a new library, and you want it to be of the greatest possible use to the public, we recommend making it free software that everyone can redistribute and change. You can do so by permitting redistribution under these terms (or, alternatively, under the terms of the ordinary General Public License).

To apply these terms, attach the following notices to the library. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the "copyright" line and a pointer to where the full notice is found.

<one line to give the library’s name and a brief idea of what it does.>

Copyright (C) <year> <name of author>

This library is free software; you can redistribute it and/or modify it under the terms of the GNU Library General Public License as published by the Free Software Foundation; either version 2 of the License, or (at your option) any later version.

This library is distributed in the hope that it will be useful, but WITHOUT ANY WARRANTY; without even the implied warranty of MERCHANTABILITY or FIT-NESS FOR A PARTICULAR PURPOSE. See the GNU Library General Public License for more details.

You should have received a copy of the GNU Library General Public License along with this library; if not, write to the Free Foundation, Inc., 59 Temple Place, Suite 330, Boston, MA 02111-1307 USA

Also add information on how to contact you by electronic and paper mail.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a "copyright disclaimer" for the library, if necessary. Here is a sample; alter the names:

Yoyodyne, Inc., hereby disclaims all copyright interest in the library ‘Frob’ (a library for tweaking knobs) written by James Random Hacker.

<signature of Ty Coon>, 1 April 1990 Ty Coon, President of Vice

That’s all there is to it!

F.2 Ispell Copyright

Copyright (c), 1983, by Pace Willisson

Copyright 1992, 1993, Geoff Kuenning, Granada Hills, CA All rights reserved.

Redistribution and use in source and binary forms, with or without modification, are permitted provided that the following conditions are met:

1. Redistributions of source code must retain the above copyright notice, this list of conditions and the following disclaimer.

2. Redistributions in binary form must reproduce the above copyright notice, this list of conditions and the following disclaimer in the documentation and/or other materials provided with the distribution.

3. All modifications to the source code must be clearly marked as such. Binary redistributions based on modified source code must be clearly marked as modified versions in the documentation and/or other materials provided with the distribution.

4. All advertising materials mentioning features or use of this software must display the following acknowledgment:

This product includes software developed by Geoff Kuenning and other unpaid contributors.

5. The name of Geoff Kuenning may not be used to endorse or promote products derived from this software without specific prior written permission.

THIS SOFTWARE IS PROVIDED BY GEOFF KUENNING AND CONTRIBUTORS “AS IS” AND ANY EXPRESS OR IMPLIED WARRANTIES, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICU-LAR PURPOSE ARE DISCLAIMED. IN NO EVENT SHALL GEOFF KUENNING OR CON-TRIBUTORS BE LIABLE FOR ANY DIRECT, INDIRECT, INCIDENTAL, SPECIAL, EXEM-PLARY, OR CONSEQUENTIAL DAMAGES (INCLUDING, BUT NOT LIMITED TO, PRO-CUREMENT OF SUBSTITUTE GOODS OR SERVICES; LOSS OF USE, DATA, OR PROFITS; OR BUSINESS INTERRUPTION) HOWEVER CAUSED AND ON ANY THEORY OF LIA-BILITY, WHETHER IN CONTRACT, STRICT LIABILITY, OR TORT (INCLUDING NEGLI-GENCE OR OTHERWISE) ARISING IN ANY WAY OUT OF THE USE OF THIS SOFTWARE, EVEN IF ADVISED OF THE POSSIBILITY OF SUCH DAMAGE.